### v1

# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

Abstract
The performance of a model trained with noisy labels is often improved by simply retraining the model with its own predicted hard labels (i.e., $1$/00 labels). Yet, a detailed theoretical characterization of this phenomenon is lacking. In this paper, we theoretically analyze retraining in a linearly separable setting with randomly corrupted labels given to us and prove that retraining can improve the population accuracy obtained by initially training with the given (noisy) labels. To the best of our knowledge, this is the first such theoretical result. Retraining finds application in improving training with label differential privacy (DP) which involves training with noisy labels. We empirically show that retraining selectively on the samples for which the predicted label matches the given label significantly improves label DP training at no extra privacy cost; we call this consensus-based retraining. For e.g., when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining.
1 Introduction
We study the simple idea of retraining an already trained model with its own predicted hard labels (i.e., $1$/00 labels and not the raw probabilities) when the given labels with which the model is initially trained are noisy. This is a simple yet effective way to boost a model’s performance in the presence of noisy labels. More formally, suppose we train a discriminative model $\mathcal{M}$ (for a classification problem) on a dataset of $n$ samples and noisy label pairs $\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j=1}^{n}$. Let $\widehat{\bm{\theta}}_{0}$ be the final learned weight/checkpoint of $\mathcal{M}$ and let $\widetilde{\mathrm{y}}_{j}=\mathcal{M}(\widehat{\bm{\theta}}_{0},{\mathbf{x}}_{j})$ be the current checkpoint’s predicted hard label for sample ${\mathbf{x}}_{j}$. Now, we propose to retrain $\mathcal{M}$ with the $\widetilde{\mathrm{y}}_{j}$’s in one of two ways.
(i) Full retraining: Retrain $\mathcal{M}$ with $\{({\mathbf{x}}_{j},\widetilde{\mathrm{y}}_{j})\}_{j=1}^{n}$, i.e., retrain $\mathcal{M}$ with the predicted labels of all the samples.
(ii) Consensus-based retraining: Define $\mathcal{S}_{\text{cons}}:=\{j\in\{1,\ldots,n\}\text{ }|\text{ }\widetilde{\mathrm{y}}_{j}=\widehat{\mathrm{y}}_{j}\}$ to be the set of samples for which the predicted label matches the given noisy label; we call this the consensus set. Retrain $\mathcal{M}$ with $\{({\mathbf{x}}_{j},\widetilde{\mathrm{y}}_{j})\}_{j\in\mathcal{S}_{\text{cons}}}$, i.e., retrain $\mathcal{M}$ with the predicted labels of only the consensus set.
Intuitively, retraining with predicted hard labels can be beneficial when the underlying classes are well-separated. In such a case, the model can potentially correctly predict the labels of many samples in the training set far away from the decision boundary which were originally incorrectly labeled and presented to it. As a result, the model’s accuracy (w.r.t. the actual labels) on the training data can be significantly higher than the accuracy of the noisy labels presented to it. Hence, retraining with predicted labels can potentially improve the model’s performance. This intuition is illustrated in Figure 1 where we consider a separable binary classification problem with noisy labels. The exact details are in Appendix A but importantly, Figures 1(a) and 1(b) correspond to versions of this problem with “large” and “small” separation, respectively. Please see the figure caption for detailed discussion but in summary, Figure 1 shows us that the success of retraining depends on the degree of separation between the classes.
The motivation for consensus-based retraining is that matching the predicted and given labels can potentially yield a smaller but much more accurate subset compared to the entire set; such a filtering effect can further improve the model’s performance. As we show in Section 5 (see Tables 3, 4 and 6), this intuition bears out in practice.
(a) Large separation: predicted labels pretty
accurate.
(b) Small separation: predicted labels not as
accurate.
Figure 1: Retraining Intuition. Samples to the right (respectively, left) of the separator (black vertical line in the middle) and colored blue (respectively, red) have actual label $+1$ (respectively, $-1$). For both classes, the incorrectly labeled samples are marked by crosses ($\times$), whereas the correctly labeled samples are marked by dots ($\circ$) of the appropriate color. The amount of label noise and the number of training samples are the same in 1(a) and 1(b). The top and bottom plots show the joint scatter plot of the training samples with the (noisy) labels given to us and the labels predicted by the model after training with the given labels, respectively. Notice that in 1(a), the model correctly predicts the labels of several samples that were given to it with the wrong label – especially, those that are far away from the separator. This is not quite the case in 1(b). This difference gets reflected in the performance on the test set after retraining. Specifically, in 1(a), retraining increases the test accuracy to $97.67\%$ from $89\%$. However, retraining yields no improvement in 1(b). So the success of retraining depends on the degree of separation between the classes.
There are plenty of ideas revolving around training a model with its own predictions, the two most common ones being self-training (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013) and self-distillation (Furlanello et al.,, 2018; Mobahi et al.,, 2020); we discuss these and important differences from retraining in Section 2. However, from a theoretical perspective, we are not aware of any work proving that retraining a model with its predicted hard labels can be beneficial in the presence of label noise in any setting. In Section 4, we derive the first theoretical result (to our knowledge) showing that full retraining with hard labels improves model accuracy.
The primary reason for our interest in retraining is that it turned out to be a simple yet effective way to improve training with label differential privacy (DP) whose goal is to safeguard the privacy of labels in a supervised ML problem by injecting label noise (see Section 3 for a formal definition). Label DP is used in scenarios where only the labels are considered sensitive, e.g., advertising, recommendation systems, etc. (Ghazi et al.,, 2021). Importantly, retraining can be applied on top of any label DP training algorithm at no extra privacy cost. Our main algorithmic contribution is empirically demonstrating the efficacy of consensus-based retraining in improving label DP training (Section 5). Three things are worth clarifying here. First, as a meta-idea, retraining is not particularly new; however, its application – especially with consensus-based filtering – as a light-weight way to improve label DP training at no extra privacy cost is new to our knowledge. Second, we are not advocating consensus-based retraining as a SOTA general-purpose algorithm for learning with noisy labels. Third, we do not view full retraining to be an algorithmic contribution; we consider it for theoretical analysis and as a baseline for consensus-based retraining.
Our main contributions can be summarized as follows:
(a) In Section 4, we consider a linearly separable binary classification problem wherein the data (feature) dimension is $d$, and we are given randomly flipped labels with the label flipping probability being $p<\frac{1}{2}$ independently for each sample. Our main result is proving that full retraining with the predicted hard labels improves the population accuracy obtained by initially training with the given labels, provided the degree of separation between the classes is large enough and the number of samples $n\in\Big{(}\Theta\Big{(}\frac{d\log d}{(1-2p)^{2}}\Big{)},\Theta\Big{(}\frac{d^{2}}{(1-2p)^{2}}\Big{)}\Big{)}$; see Remark 4 for details. Our results also show that retraining becomes more beneficial as the amount of label noise (i.e., $p$) increases (Remark 4) or as the degree of separation between the classes increases (Remark 5). To the best our knowledge, these are the first theoretical results quantifying the benefits of retraining with predicted hard labels in the presence of label noise.
(b) In Section 5, we show the promise of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) as a simple way to improve the performance of any label DP algorithm, at no extra privacy cost. For e.g., when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining (see Table 2). The corresponding improvement for a small BERT model trained on AG News Subset (a news classification dataset) with $\epsilon=0.5$ label DP is $11.7\%$ (see Table 5).
2 Related Work
Self-Training (ST). Retraining is similar in spirit to ST (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013; Sohn et al.,, 2020) which is the process of progressively training a model with its own predicted hard labels in the semi-supervised setting. Our main focus is the fully supervised setting; [^1] It is worth mentioning here that we analyze retraining in a semi-supervised setting due to certain technical complications (because of independence not holding) in the supervised setting. But we believe that the crux of our theoretical insights should carry over to the supervised setting. in fact, our proposed consensus-based retraining scheme crucially relies on the given labels for sample selection. This is different from ST (in the semi-supervised setting) which typically selects samples based on the model’s confidence and hence, we call our algorithmic idea of interest retraining to distinguish it from ST. In fact, we show that our consensus-based sample selection strategy leads to better performance than confidence-based sample selection in Appendix I. There is a vast body of work on ST and related ideas; see Amini et al., (2022) for a survey. On the theoretical side also, there are several papers showing and quantifying different kinds of benefits of ST and related ideas (Carmon et al.,, 2019; Raghunathan et al.,, 2020; Kumar et al.,, 2020; Chen et al.,, 2020; Oymak and Gulcu,, 2020; Wei et al.,, 2020; Zhang et al.,, 2022). But none of these works characterize the pros/cons of ST or any related algorithm in the presence of noisy labels. In contrast, we show that retraining can provably improve accuracy in the presence of label noise in Section 4. Empirically, ST-based ideas have been proposed to improve learning with noisy labels (Reed et al.,, 2014; Tanaka et al.,, 2018; Han et al.,, 2019; Nguyen et al.,, 2019; Li et al.,, 2020; Goel et al.,, 2022); but these works do not have rigorous theory.
Self-Distillation (SD). Retraining is also similar in principle to SD (Furlanello et al.,, 2018; Mobahi et al.,, 2020); the major difference is that soft labels (i.e., predicted raw probabilities) are used in SD, whereas we use hard labels in retraining. Specifically, in SD, a teacher model is first trained with provided hard labels and then its predicted soft labels are used to train a student model with the same architecture as the teacher. SD is usually employed with a temperature parameter (Hinton et al.,, 2015) to force the teacher and student models to be different; we do not have any such parameter in retraining as it uses hard labels. SD is known to ameliorate learning in the presence of noisy labels (Li et al.,, 2017) and this has been theoretically analyzed by Dong et al., (2019); Das and Sanghavi, (2023). Dong et al., (2019) propose their own SD algorithm that uses dynamically updated soft labels and provide some complicated conditions of when their algorithm can learn the correct labels in the presence of noisy labels. In contrast, we analyze retraining with fixed hard labels. Das and Sanghavi, (2023) analyze the standard SD algorithm in the presence of noisy labels with fixed soft labels but their analysis in the classification setting requires some strong assumptions such as access to the population, feature maps of all points in the same class having the same inner product, etc. We do not require such strong assumptions in this paper (in fact, we present sample complexity bounds). Moreover, Das and Sanghavi, (2023) have extra $\ell_{2}$-regularization in their objective function to force the teacher and student models to be different. We do not apply any extra regularization for retraining.
Label Differential Privacy (DP). Label DP (described in detail in Section 3) is a relaxation of full-data DP wherein the privacy of only the labels (and not the features) is safeguarded (Chaudhuri and Hsu,, 2011; Beimel et al.,, 2013; Wang and Xu,, 2019; Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021; Ghazi et al.,, 2022; Badanidiyuru et al.,, 2023). In this work, we are not trying to propose a SOTA label DP algorithm (with an ingenious noise-injection scheme); instead, we advocate retraining as a simple post-processing step that can be applied on top of any label DP algorithm (regardless of the noise-injection scheme) to improve its performance, at no extra privacy cost. Similar to our goal, Tang et al., (2022) apply techniques from unsupervised and semi-supervised learning to improve label DP training. In particular, one of their steps involves keeping the given noisy label of a sample only if it matches a pseudo-label generated by unsupervised learning. This is similar in spirit to our consensus-based retraining scheme but a crucial difference is that we do not perform any unsupervised learning; we show that matching the given noisy label to the model’s own predicted label is itself pretty effective. Further, unlike work, Tang et al., (2022) do not have any rigorous theory.
3 Preliminaries
Notation. We use $\Theta(\cdot)$ and $\Omega(\cdot)$ to denote the standard big-$\Theta$ and big-$\Omega$ notations. Specifically, $f(d)=\Theta(g(d))$ if there are absolute constants $c_{1},c_{2}>0$ and a natural number $d_{0}$ such that $c_{1}g(d)\leq f(d)\leq c_{2}g(d)$ for all $d\geq d_{0}$. In addition, $f(d)=\Omega(g(d))$ if there is an absolute constant $c$ and a natural number $d_{0}$ such that $f(d)\geq cg(d)$ for all $d\geq d_{0}$. Throughout the paper, we sometimes use the big-$\Theta$ notation to absorb different absolute constants. For any positive integer $m\geq 1$, we denote the set $\{1,\ldots,m\}$ by $[m]$. For any vector ${\mathbf{v}}$, we denote its $i^{\text{th}}$ coordinate by $\mathrm{v}^{(i)}$. Let ${\mathbf{e}}_{i}$ denote the $i^{\text{th}}$ canonical vector, i.e., the vector of all zeros except a one in the $i^{\text{th}}$ coordinate. We denote the $\ell_{2}$ norm of a vector ${\mathbf{v}}$ by $\|{\mathbf{v}}\|$, and the operator norm of a matrix ${\mathbf{M}}$ by $\|{\mathbf{M}}\|$. The unit $d$-dimensional sphere (i.e., the set of $d$-dimensional vectors with unit norm) is denoted by $S^{d-1}$. A random variable ${\mathrm{X}}$ is said to be sub-Gaussian with parameter $\sigma^{2}$ (abbreviated as SG($\sigma^{2}$)) if for all $t\in\mathbb{R}$, it holds that $\mathbb{E}\big{[}e^{t({\mathrm{X}}-\mathbb{E}[{\mathrm{X}}])}\big{]}\leq\exp\big{(}\frac{t^{2}\sigma^{2}}{2}\big{)}$. We denote the CDF and complementary CDF (CCDF) of a standard normal variable (i.e., distributed as $\mathcal{N}(0,1)$) by $\Phi(.)$ and $\Phi^{\text{c}}(.)$, respectively.
Definition 1 (Label Differential Privacy (DP)).
A randomized algorithm $\mathcal{A}$ taking as input a dataset and with range $\mathcal{R}$ is said to be $\epsilon$-labelDP if for any two datasets ${\mathrm{D}}$ and ${\mathrm{D}}^{\prime}$ differing in the label of a single example and for any $\mathcal{S}\subseteq\mathcal{R}$, it holds that $\mathbb{P}\big{(}\mathcal{A}({\mathrm{D}})\in\mathcal{S}\big{)}\leq e^{\epsilon}\mathbb{P}\big{(}\mathcal{A}({\mathrm{D}}^{\prime})\in\mathcal{S}\big{)}$.
Label DP training involves injecting noise into the labels and then training with these noisy labels. The simplest way of injecting label noise to ensure label DP is randomized response (RR) introduced by Warner, (1965). Specifically, suppose we require $\epsilon$-labelDP for a problem with $C$ classes, then the distribution of the output $\widehat{\mathrm{y}}$ of RR when the true label is ${\mathrm{y}}$ is as follows:
$\mathbb{P}\big{(}\hat{\mathrm{y}}={\mathrm{z}}\big{)}=\begin{cases}\frac{e^{\epsilon}}{e^{\epsilon}+C-1}\quad\text{ for }{\mathrm{z}}={\mathrm{y}},\\ \frac{1}{e^{\epsilon}+C-1}\quad\text{ otherwise}.\end{cases}$
Our label noise model in Section 4 (eq. 1) is actually RR for 2 classes. Based on vanilla RR, more sophisticated ways to inject label noise for better performance under label DP have been proposed (Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021). Our empirical results in Section 5 are with RR and the method of Ghazi et al., (2021).
4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
Here we will analyze full retraining (as introduced in Section 1) for a linear setting with noisy labels. Since full retraining is the only kind of retraining we consider in this section, we will omit the word “full” subsequently in this section.
[…]
We are given $n>d$ labeled examples $\mathcal{X}_{L}:=\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}\}$ and $n$ unlabeled examples $\mathcal{X}_{U}:=\{{\mathbf{x}}^{\prime}_{1},\ldots,{\mathbf{x}}^{\prime}_{n}\}$ drawn i.i.d. from $\mathcal{D}$. For $j\in[n]$, let $\mathrm{y}_{j}=\text{sign}\big{(}\langle{\mathbf{x}}_{j},\bm{\theta}^{\ast}\rangle\big{)}$ be the ground truth label of ${\mathbf{x}}_{j}$. We have access to a noisy label source which gives us a randomly corrupted version of $\mathrm{y}_{j}$, namely $\widehat{\mathrm{y}}_{j}$, instead of $\mathrm{y}_{j}$. Specifically, for some $p<\frac{1}{2}$, we have: [^5] In the context of $\epsilon$-labelDP with randomized response, $p=\frac{1}{1+e^{\epsilon}}$.
$\widehat{\mathrm{y}}_{j}=\begin{cases}\mathrm{y}_{j}\text{ w.p. }1-p\\ -\mathrm{y}_{j}\text{ w.p. }p.\end{cases}$ (1)
Let $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$ denote the labeled part of the training data. We use $\mathcal{T}$ for vanilla training. Then, we label $\mathcal{X}_{U}$ using the classifier learned with vanilla training and use this for retraining. Ideally, we would have liked to analyze retraining over $\mathcal{X}_{L}$ and not have $\mathcal{X}_{U}$ at all, but this is complicated and tedious due to independence not holding. In order to circumvent this extra layer of complication in the analysis with $\mathcal{X}_{L}$, we analyze retraining over $\mathcal{X}_{U}$ (wherein independence obviates the said complication). We believe that the main insights should not change much.
4.1 Vanilla Training
Here we train with the given noisy labels as is. Our objective function with the squared loss is $\frac{1}{2n}\sum_{j=1}^{n}\big{(}\widehat{\mathrm{y}}_{j}-\bm{\theta}^{\top}{\mathbf{x}}_{j}\big{)}^{2}$ and its minimizer is $\bar{\bm{\theta}}_{0}=\Big{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}\Big{)}^{-1}\Big{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Big{)}$. But instead of considering $\bar{\bm{\theta}}_{0}$ as our learned separator, we will focus on
$\widehat{\bm{\theta}}_{0}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}.$ (2)
Observe that the empirical covariance matrix $\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}$ in $\bar{\bm{\theta}}_{0}$ is replaced by its expected value, viz., $\sigma^{2}{\mathbf{I}}_{d}$ in $\widehat{\bm{\theta}}_{0}$. The reason for doing this is that the analysis with $\bar{\bm{\theta}}_{0}$ eventually reduces down to analyzing with $\widehat{\bm{\theta}}_{0}$ by performing some tedious math first (due to the empirical covariance matrix) which does not yield any meaningful insights. For completeness, we discuss how the analysis can be done with $\bar{\bm{\theta}}_{0}$ in Appendix G. Our predicted label for a sample ${\mathbf{x}}$ with $\widehat{\bm{\theta}}_{0}$ is $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$ (instead of just $\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle$ which would have been the case in a regression problem).
Theorem 1 (Vanilla Training: Probability of Predicted Label = True Label).
Consider a sample ${\mathbf{x}}\notin\mathcal{X}_{L}$. Let $\beta({\mathbf{x}})$ denote the angle between ${\mathbf{x}}$ and $\bm{\theta}^{\ast}$. Then our predicted label for ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$ is equal to the ground truth label of ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ with a probability (over the randomness of $\mathcal{T}$) of at least $\alpha({\mathbf{x}}):=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ and at most $\widehat{\alpha}({\mathbf{x}}):=1-\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}$.
The proof of Theorem 1 is in Appendix B. Observe that $\alpha({\mathbf{x}})$ and $\widehat{\alpha}({\mathbf{x}})$ have essentially the same dependence on $n$, $p$ and $\mu$; so, this dependence is tight. Also, notice that the learned classifier $\widehat{\bm{\theta}}_{0}$ is more likely to be wrong on the samples that are less aligned with (or closer to orthogonal to) the ground truth separator, i.e., $\bm{\theta}^{\ast}$. We can view $\widehat{\bm{\theta}}_{0}$ as a noisy label provider (a.k.a. pseudo-labeler) where the degree of label noise is non-uniform or sample-dependent unlike the original noisy source used to learn $\widehat{\bm{\theta}}_{0}$. Specifically, for a sample ${\mathbf{x}}$ with true label $\mathrm{y}=\text{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ and predicted label $\widetilde{\mathrm{y}}=\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}$, we have:
[…]
We will now provide lower and upper bounds on the population accuracy of the classifier $\widehat{\bm{\theta}}_{0}$; we denote this by $\text{acc}(\widehat{\bm{\theta}}_{0})$ and it is defined as:
$\text{acc}(\widehat{\bm{\theta}}_{0}):=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}\Big{)}.$
Theorem 2 (Vanilla Training: Population Accuracy).
We have:
$\textup{acc}(\widehat{\bm{\theta}}_{0})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Bigg{)},$
and
$\textup{acc}(\widehat{\bm{\theta}}_{0})\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$
The proof of Theorem 2 is in Appendix C.
Remark 1 (Tightness of Accuracy Bounds).
Suppose $\gamma=\Theta(\mu)$. When $n\leq\Theta\Big{(}\frac{d^{2}}{(1-2p)^{2}\mu^{4}}\Big{)}$, the lower and upper bounds for $\textup{acc}(\widehat{\bm{\theta}}_{0})$ in Theorem 2 essentially match (modulo constant factors).
Based on Theorem 2, we have the following corollary.
Corollary 1 (Vanilla Training: Sample Complexity).
For any $\delta>(1+\Theta(1))\exp(-\Theta({d}))$, $n=\Omega\Big{(}\frac{\log{1}/{\delta}}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}\Big{)}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-\delta$. Specifically, when $p>(1+\Theta(1))\exp(-\Theta({d}))$, $n=\Omega\Big{(}\frac{\log{1}/{p}}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}\Big{)}$ ensures that $\textup{acc}(\widehat{\bm{\theta}}_{0})>1-p$, i.e., our learned classifier $\widehat{\bm{\theta}}_{0}$ has better accuracy than the source providing noisy labels (used to learn $\widehat{\bm{\theta}}_{0}$).
We will now present an information-theoretic lower bound on the sample complexity of any classifier to argue the optimality of the result in Corollary 1 with respect to the dependence on $d$ and $p$.
Theorem 3 (Information-Theoretic Lower Bound on Sample Complexity).
With a slight generalization of notation, let $\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})$ denote the accuracy of the classifier $\widehat{\bm{\theta}}$, when the ground truth model is $\bm{\theta}^{*}$. For any classifier $\widehat{\bm{\theta}}$ learned from $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$, in order to achieve $\inf_{\bm{\theta}^{\ast}\in S^{d-1}}\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})\geq 1-\delta$, where $\delta<\frac{1}{5}$, $n=\Omega\left(\frac{(1-5\delta)}{(1-2p)^{2}}d\right)$ is necessary in our problem setting.
It is worth mentioning that there is a similar lower bound in Gentile and Helmbold, (1998) for a different classification setting. In contrast, Theorem 3 is tailored to our setting and moreover, the proof technique is also different. Specifically, for the proof of Theorem 3, we follow a standard technique in proving minimax lower bounds which is to reduce the problem of interest to an appropriate multi-way hypothesis testing problem; this is accompanied by the application of the conditional version of Fano’s inequality and some ideas from high-dimensional geometry. This proof is in Appendix D.
Remark 2 (Optimality of Sample Complexity).
Note that the dependence of the sample complexity on $d$ and $p$ in Corollary 1 matches that of the lower bound in Theorem 3. Thus, our sample complexity bound in Corollary 1 is optimal with respect to $d$ and $p$.
Remark 3 (Effect of Degree of Separation).
As the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy lower bound in Theorem 2 also decreases and the sample complexity required to outperform the noisy label source in Corollary 1 increases. This is consistent with our intuition that a classification task should become harder as the degree of separation reduces; we also saw this in Figure 1.
4.2 Retraining
We first label the unlabeled set $\mathcal{X}_{U}:=\{{\mathbf{x}}^{\prime}_{1},\ldots,{\mathbf{x}}^{\prime}_{n}\}$ using the learned classifier $\widehat{\bm{\theta}}_{0}$. Let us denote the predicted label for ${\mathbf{x}}^{\prime}_{j}$ with $\widehat{\bm{\theta}}_{0}$ by $\widetilde{\mathrm{y}}_{j}^{\prime}=\textup{sign}\big{(}\langle{\mathbf{x}}^{\prime}_{j},\widehat{\bm{\theta}}_{0}\rangle\big{)}$. Also, let $\mathcal{T}_{2}:=\{({\mathbf{x}}^{\prime}_{j},\widetilde{\mathrm{y}}_{j}^{\prime})\}_{j\in[n]}$; we use $\mathcal{T}_{2}$ for retraining. Just like Section 4.1 (using the squared loss), our separator of interest with $\mathcal{T}_{2}$ is:
$\widehat{\bm{\theta}}_{1}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widetilde{\mathrm{y}}_{j}^{\prime}{\mathbf{x}}_{j}^{\prime}\Bigg{)}.$ (3)
Theorem 4 (Retraining: Probability of Predicted Label = True Label).
Consider a sample ${\mathbf{x}}\notin\mathcal{X}_{L}\cup\mathcal{X}_{U}$. Let $\beta({\mathbf{x}})$ denote the angle between ${\mathbf{x}}$ and $\bm{\theta}^{\ast}$. Also, let $p_{2}:=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}$. As long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{16p_{2}}{\mu}$, our predicted label for ${\mathbf{x}}$, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}$ is equal to its ground truth label, viz., $\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}$ with a probability (over the randomness of $\mathcal{T}$ and $\mathcal{T}_{2}$) of at least $\alpha_{1}({\mathbf{x}}):=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$.
The proof of Theorem 4 is in Appendix E; it involves non-trivial moment generating function (MGF) computations. The analysis is especially challenging because as we discussed after Theorem 1, the classifier learned with vanilla training, i.e., $\widehat{\bm{\theta}}_{0}$, is a non-uniform noisy label provider. Also, just like the classifier $\widehat{\bm{\theta}}_{0}$, the classifier $\widehat{\bm{\theta}}_{1}$ learned with retraining is more likely to be wrong on the samples that are less aligned with the ground truth separator $\bm{\theta}^{\ast}$.
We will now provide a lower bound on the population accuracy of the classifier $\widehat{\bm{\theta}}_{1}$; we denote this by $\text{acc}(\widehat{\bm{\theta}}_{1})$ and it is defined as:
$\text{acc}(\widehat{\bm{\theta}}_{1}):=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}\langle{\mathbf{x}},\bm{\theta}^{\ast}\rangle\big{)}\Big{)}.$
Theorem 5 (Retraining: Population Accuracy).
Suppose $p>(1+\Theta(1))\exp(-\Theta({d}))$ and $\gamma\geq\Theta\Big{(}\frac{\sqrt{d}\exp(-\Theta(d))}{\mu}\Big{)}$. Let $n\geq\Theta\Big{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\max\Big{(}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)},\log\frac{1}{p}\Big{)}\Big{)}$. Then:
$\textup{acc}(\widehat{\bm{\theta}}_{1})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)},$
with $p_{2}=\exp\big{(}-\Theta(d)\big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}\leq\min\Big{(}p,\Theta\big{(}\frac{\mu\gamma}{\sqrt{d}}\big{)}\Big{)}$.
The proof of Theorem 5 is in Appendix F. Note that the sample complexity required for Theorem 5 to hold is more than the lower bound in Theorem 3 (specifically, there is an extra $\log\big{(}\frac{d}{\mu^{2}\gamma^{2}}\big{)}$ term in Theorem 5). So Theorem 3 is never violated by Theorem 5.
We will now compare the accuracies of vanilla training and retraining.
Remark 4 (When Does Retraining Improve Accuracy?).
Suppose we are in the non-trivial noise regime where $p$ is a constant (e.g., $\frac{2}{5}$). Moreover, suppose $\gamma=\Theta(\mu)$. Then, as per Theorem 5, when $n\geq\Theta\Big{(}\frac{1}{(1-2p)^{2}}\big{(}\frac{d}{\mu^{4}}\big{)}\log\big{(}\frac{d}{\mu^{4}}\big{)}\Big{)}$, [^6] This follows because $\gamma=\Theta(\mu)$ and $p$ is a constant. the accuracy of retraining is $\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Big{(}-\frac{n\mu^{4}}{\Theta(d)}\Big{)}$. [^7] This holds because $\gamma=\Theta(\mu)$ and $p_{2}\leq\Theta\Big{(}\frac{\mu^{2}}{\sqrt{d}}\Big{)}$, since $p$ is a constant. This can be further lower-bounded by:
$\geq 1-\Theta(1)\exp\Bigg{(}-\min\Bigg{(}\Theta(d),\frac{n\mu^{4}}{\Theta(d)}\Bigg{)}\Bigg{)}.$ (4)
In comparison, the accuracy of vanilla training as per Theorem 2 is:
$\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (5)
From eq. 4 and eq. 5 and as per the lower bound on $n$ stated before, when
$n\in\Bigg{(}\Theta\Bigg{(}\frac{(d/\mu^{4})\log\big{(}{{d}/{\mu^{4}}}\big{)}}{(1-2p)^{2}}\Bigg{)},\Theta\Bigg{(}\frac{(d^{2}/\mu^{4})}{(1-2p)^{2}}\Bigg{)}\Bigg{)}$ (6)
and $p$ is sufficiently close to $\frac{1}{2}$, the lower bound on the accuracy of retraining (in eq. 4) is greater than the upper bound on the accuracy of vanilla training (in eq. 5). This improvement becomes increasingly significant as $p$ approaches $\frac{1}{2}$.
Regarding $n$ in Remark 4. We expect that the minimum value of $n$ needed for retraining to be beneficial should be at least the sample complexity needed for the classifier obtained with vanilla training to outperform the noisy label source used to train it. We computed the latter in Corollary 1; compared to this value, the minimum value of $n$ required for Remark 4 to hold (in eq. 6) has an extra mild factor of $\log\big{(}{\frac{d}{\mu^{4}}}\big{)}$. [^8] Recall that this is when $\gamma=\Theta(\mu)$ and when $p$ is a constant. On the other hand, the upper bound on $n$ in eq. 6 of Remark 4 could be an artifact of our analysis; investigating this is left for future work. [^9] More specifically, this upper bound on $n$ arises as a result of the $\exp\big{(}-\Theta(d)\big{)}$ term in the accuracy bound of retraining. If this can be removed, then the upper bound will not exist.
Remark 5 (Effect of Degree of Separation).
Similar to vanilla training (Remark 3), as the parameter quantifying the degree of separation $\gamma$ decreases, the accuracy lower bound of retraining in Theorem 5 also decreases [^10] Note that $p_{2}$ also depends on $\gamma$ and it is a decreasing function of $\gamma$. and the sample complexity required for this bound to hold increases. Moreover, in Remark 4, we considered the regime of $\gamma=\Theta(\mu)$, i.e., $\gamma$ is “large” enough relative to the mean $\mu$. This tells us that retraining is more beneficial when the degree of separation is large.
5 Improving Label DP Training with Retraining (RT)
Motivated by our theoretical results in Section 4 which show that retraining (abbreviated as RT henceforth) can improve accuracy in the presence of label noise, we propose to apply our proposals in Section 1, viz., full RT and more importantly, consensus-based RT to improve label DP training (because it involves noisy labels). Note that this can be done on top of any label DP mechanism and that too at no additional privacy cost (both the predicted labels and originally provided noisy labels are private). We empirically evaluate full and consensus-based RT on three classification datasets (available on TensorFlow) trained with label DP. These include two vision datasets, namely CIFAR-10 and CIFAR-100, and one language dataset, namely AG News Subset (Zhang et al.,, 2015). All the empirical results are averaged over three different runs. We only provide important experimental details here; the other details can be found in Appendix H.
CIFAR-10/100. We train a ResNet-18 model on CIFAR-10 and CIFAR-100 with label DP. Label DP training is done with the prior-based method of Ghazi et al., (2021) – specifically, Alg. 3 with two stages. Our training set consists of 45k examples and we assume access to a validation set with clean labels consisting of 5k examples which we use for deciding when to stop training, setting hyper-parameters, etc. [^11] In practice, we do not need full access to the validation set. Instead, the validation set can be stored by a secure agent which returns us a private version of the validation accuracy and this will not be too far off from the true validation accuracy when the validation set is large enough. For CIFAR-10 and CIFAR-100 with three different values of $\epsilon$, we list the test accuracies of the baseline (i.e., the method of Ghazi et al., (2021)), full RT and consensus-based RT in Tables 1 and 2, respectively. Notice that consensus-based RT is the clear winner. Also, for the three values of $\epsilon$ in Table 1 (CIFAR-10), the size of the consensus set (used in consensus-based RT) is $\sim 31\%$, $55\%$ and $76\%$, respectively, of the entire training set. The corresponding numbers for Table 2 (CIFAR-100) are $\sim 11\%$, $34\%$ and $56\%$, respectively. So for small $\epsilon$ (high label noise), consensus-based RT comprehensively outperforms full RT and baseline with a small fraction of the training set. Further, in Tables 3 and 4, we list the accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set for CIFAR-10 and CIFAR-100. Please see the table captions for detailed discussion but to summarize, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than the full dataset.
Table 1: CIFAR-10. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT is better than full RT which is better than the baseline.
$\epsilon$ Baseline Full RT Consensus-based RT
1 $57.78\pm 1.13$ $60.07\pm 0.63$ $\textbf{63.84}\pm 0.56$
2 $79.06\pm 0.59$ $81.34\pm 0.40$ $\textbf{83.31}\pm 0.28$
3 $85.18\pm 0.50$ $86.67\pm 0.28$ $\textbf{87.67}\pm 0.28$
Table 2: CIFAR-100. Test set accuracies (mean $\pm$ standard deviation). Overall, consensus-based RT is significantly better than full RT which is somewhat better than the baseline.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $23.53\pm 1.01$ $24.42\pm 1.22$ $\textbf{29.98}\pm 1.11$
4 $44.53\pm 0.81$ $46.99\pm 0.66$ $\textbf{51.30}\pm 0.98$
5 $55.75\pm 0.36$ $56.98\pm 0.43$ $\textbf{59.47}\pm 0.26$
Table 3: CIFAR-10. Accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set. Note that the accuracy of predicted labels over the consensus set $\gg$ accuracy of predicted labels over the entire dataset $\gg$ accuracy of given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is smaller than the full dataset ($\sim$ $31\%$, $55\%$ and $76\%$ of the full dataset for $\epsilon=1,2$ and $3$, respectively).
$\epsilon$ Acc. of predicted labels on full dataset Acc. of given labels on full dataset Acc. of predicted labels on consensus set
1 $59.30\pm 0.74$ $32.61\pm 0.74$ $\textbf{76.17}\pm 0.15$
2 $81.62\pm 0.18$ $57.11\pm 0.05$ $\textbf{92.65}\pm 0.22$
3 $89.28\pm 0.35$ $76.73\pm 0.12$ $\textbf{95.94}\pm 0.23$
Table 4: CIFAR-100. Accuracies of predicted labels and given labels over the entire (training) dataset and accuracies of predicted (= given) labels over the consensus set. Here the accuracy of predicted labels over the consensus set $\gg$ accuracy of predicted labels over the entire dataset $\approx$ accuracy of given labels over the entire dataset. This gives us an idea of why consensus-based RT is much better than full RT and baseline, even though the consensus set is much smaller than the full dataset ($\sim$ $11\%$, $34\%$ and $56\%$ of the full dataset for $\epsilon=3,4$ and $5$, respectively). But unlike CIFAR-10 (Table 3), here the accuracy of predicted labels over the entire dataset is not too much better than the given labels. Thus, full RT is only somewhat better than the baseline for CIFAR-100 (see Table 2).
$\epsilon$ Acc. of predicted labels on full dataset Acc. of given labels on full dataset Acc. of predicted labels on consensus set
3 $24.90\pm 0.92$ $22.35\pm 0.41$ $\textbf{76.09}\pm 0.85$
4 $50.85\pm 0.82$ $46.32\pm 0.34$ $\textbf{91.59}\pm 1.24$
5 $66.51\pm 0.02$ $68.09\pm 0.33$ $\textbf{94.83}\pm 0.15$
AG News Subset (https://www.tensorflow.org/datasets/catalog/ag_news_subset). This is a news article classification dataset consisting of 4 categories – world, sports, business or sci/tech. We reserve 10% of the given training set for validation and we use the rest for training with label DP. Just like the CIFAR experiments, we assume that the validation set comes with clean labels. We use the small BERT model available in TensorFlow and the BERT English uncased preprocessor; links to both of these are in Appendix H. We pool the output of the BERT encoder, add a dropout layer with probability = 0.2, followed by a softmax layer. We fine-tune the full model. Here, label DP training is done with randomized response. We list the test accuracies of the baseline (i.e., randomized response), full RT and consensus-based RT in Table 5 for three different values of $\epsilon$. Even here, consensus-based RT is the clear winner. For the three values of $\epsilon$ in Table 5, the size of the consensus set (used in consensus-based RT) is $\sim 28\%$, $32\%$ and $38\%$, respectively, of the entire training set. So here, consensus-based RT appreciably outperforms full RT and baseline with less than two-fifths of the entire training set. Finally, in Table 6, we list the accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted (= given) labels over the consensus set. Even here, the accuracy of predicted labels over the consensus set is significantly more than the accuracy of predicted and given labels over the entire dataset. This explains why consensus-based RT performs the best, even though the consensus set is much smaller than the full dataset.
Table 5: AG News Subset. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT is better than full RT which is better than the baseline.
$\epsilon$ Baseline Full RT Consensus-based RT
0.3 $54.54\pm 0.97$ $60.03\pm 2.90$ $\bm{65.91}\pm 1.93$
0.5 $69.21\pm 0.31$ $75.63\pm 1.08$ $\bm{80.95}\pm 1.47$
0.8 $79.10\pm 1.43$ $82.19\pm 1.54$ $\bm{84.26}\pm 1.03$
Table 6: AG News Subset. Accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted (= given) labels over the consensus set. Conclusions are the same as Table 3.
$\epsilon$ Acc. of predicted labels on full dataset Acc. of given labels on full dataset Acc. of predicted labels on consensus set
0.3 $53.20\pm 2.82$ $32.52\pm 2.05$ $\bm{61.81}\pm 2.66$
0.5 $66.78\pm 1.31$ $35.5\pm 0.14$ $\bm{76.48}\pm 0.93$
0.8 $79.98\pm 0.80$ $42.53\pm 0.13$ $\bm{89.59}\pm 0.43$
So in summary, consensus-based RT significantly improves model accuracy. In Appendix I, we show that consensus-based RT also outperforms retraining on samples for which the model is the most confident; this is similar to self-training’s method of sample selection in the semi-supervised setting. Furthermore, going beyond label DP, we show that consensus-based RT is also beneficial in the presence of human annotation errors which can be thought of as “real” label noise in Appendix J. Also, the test accuracies without label DP for all the above datasets are listed in Appendix H.
6 Conclusion and Limitations
In this work, we provided the first theoretical result showing retraining with hard labels can provably increase model accuracy in the presence of label noise. We also showed the efficacy of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) in improving label DP training at no extra privacy cost.
We will conclude by discussing some limitations of our work which also pave the way for some future directions of work. Our retraining analysis is done on an unlabeled dataset different from the initial labeled dataset used for vanilla training (due to certain technical complications). In the future, we would like to analyze retraining on the same labeled dataset. We did not investigate the tightness of the results in Theorem 5 and Remark 4; investigating this is left for future work. Moreover, we have only analyzed full retraining in this work. Given that consensus-based retraining worked very well empirically, we would like to analyze it theoretically later. Another potential extension of our work is to analyze retraining under non-uniform label noise models. Our experiments in this paper are not on very large-scale models or datasets. We hope to test our ideas on larger scale problems in the future.
Appendix
Contents:
Appendix A Problem Setting of Figure 1
The setting is exactly the same as the problem setting in Section 4 with $\bm{\theta}^{*}={\mathbf{e}}_{1}$, =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$, $\sigma=1$, $d=50$ and $p=0.4$, except for the distribution of ${\mathrm{a}}^{(1)}$ which is equal to ${\mathrm{x}}^{(1)}$ as $\bm{\theta}^{*}={\mathbf{e}}_{1}$. Specifically for some $\Delta>0$, ${\mathrm{x}}^{(1)}$ is drawn from $\text{Unif}[\Delta,5\Delta]$ with probability $\frac{1}{2}$ or from $\text{Unif}[-5\Delta,-\Delta]$ otherwise ($\text{Unif}[a,b]$ denotes the uniform distribution over the interval $[a,b]$). Note that this distribution is similar in principle to the one assumed in Section 4 (where we did not assume ${\mathrm{a}}^{(1)}$ to have any specific distribution). In Figure 1(a), $\Delta=0.5$ (large separation) and in Figure 1(b), $\Delta=0.3$ (small separation). The number of training samples in each case is 300 and the retraining is done on the same training set on which the model is initially trained. The learned classifiers from vanilla training and retraining are the same as in Section 4 (i.e., eq. 2 and eq. 3, respectively). Finally, the test accuracy of both vanilla training and retraining in Figure 1(b) is $68\%$.
Appendix B Proof of Theorem 1
Proof.
Recall the problem setting at the beginning of Section 4. Without loss of generality, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ (i.e., the canonical basis). In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$. Let ${\mathrm{y}}=\text{sign}({\mathrm{x}}^{(1)})$ be the ground truth label of ${\mathbf{x}}$. Note that proving Theorem 1 is equivalent to proving the following statement. For a sample ${\mathbf{x}}\notin\mathcal{X}_{L}$, the following holds with a probability of at least $\alpha({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ and at most $\widehat{\alpha}({\mathbf{x}})=1-\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}$:
$\begin{cases}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\geq 0\text{ when }{\mathrm{y}}=1\text{ i.e., }{\mathrm{x}}^{(1)}\geq 0,\\ \langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle<0\text{ when }{\mathrm{y}}=-1\text{ i.e., }{\mathrm{x}}^{(1)}<0.\end{cases}$ (7)
(a) Lower bound $\alpha({\mathbf{x}})$. First note that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ in our choice of basis for the proof. We will prove the result for the case of ${\mathrm{y}}=1$ or equivalently ${\mathrm{x}}^{(1)}\geq 0$; the result for the other case (i.e., ${\mathrm{y}}=-1$ or equivalently ${\mathrm{x}}^{(1)}<0$) can be derived similarly.
We have:
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle=\frac{1}{\sigma^{2}}\Bigg{\{}\frac{1}{n}\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle\Bigg{\}}.$ (8)
Since $\frac{1}{n\sigma^{2}}$ is a constant independent of the training set $\mathcal{T}$, we will analyze
$\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle.$ (9)
Using the Chernoff bound, we have for any $r<0$:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $=\mathbb{P}\big{(}\exp\big{(}r\varepsilon\big{)}\geq 1\big{)}\leq\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}.$ (10)
For $k\in[n]$, let $\varepsilon_{k}:=\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle$; thus, $\varepsilon=\sum_{k\in[n]}\varepsilon_{k}$. For any $r\in\mathbb{R}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r\sum_{k\in[n]}\varepsilon_{k}\big{)}\Big{]}=\mathbb{E}\Big{[}\prod_{k\in[n]}\exp\big{(}r\varepsilon_{k}\big{)}\Big{]}=\Big{(}\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}\big{)}\Big{]}\Big{)}^{n},$ (11)
where the last step follows because of the i.i.d. nature of the data and label noise. Now:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}\big{)}\Big{]}$ $=\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}.$ (12)
Using the definition of $\widehat{\mathrm{y}}_{1}$ from eq. 1, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $=\Bigg{(}(1-p)\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}+p\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}\Bigg{)}.$ (13)
So now we need to evaluate $\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ and $\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$. Recalling that ${\mathrm{y}}_{1}=\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Bigg{[}\exp\Bigg{(}r\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}\Bigg{(}{\mathrm{x}}_{1}^{(1)}{\mathrm{x}}^{(1)}+\sum_{j=2}^{d}{\mathrm{x}}_{1}^{(j)}{\mathrm{x}}^{(j)}\Bigg{)}\Bigg{)}\Bigg{]},$ (14)
Simplifying the above by using the independence of ${\mathrm{x}}_{1}^{(j)}$’s and recalling that each ${\mathrm{x}}_{1}^{(j)}\sim\mathcal{N}(0,\sigma^{2})$ for $j\in\{2,\ldots,d\}$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)},{\mathrm{x}}_{1}^{(2)},\ldots,{\mathrm{x}}_{1}^{(d)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\exp\Big{(}r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}{\mathrm{x}}_{1}^{(j)}\Big{)}\Bigg{]}$ (15)
$=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\mathbb{E}_{{\mathrm{x}}_{1}^{(j)}}\Big{[}\exp\Big{(}r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}{\mathrm{x}}_{1}^{(j)}\Big{)}\Big{]}\Bigg{]}$ (16)
$=\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\prod_{j=2}^{d}\exp\Bigg{(}\frac{\big{(}\sigma r{\mathrm{x}}^{(j)}\text{sign}\big{(}{\mathrm{x}}_{1}^{(1)}\big{)}\big{)}^{2}}{2}\Bigg{)}\Bigg{]}$ (17)
$={\mathbb{E}_{{\mathrm{x}}_{1}^{(1)}}\Bigg{[}\exp\Big{(}r{\mathrm{x}}^{(1)}\big{|}{\mathrm{x}}_{1}^{(1)}\big{|}\Big{)}\Bigg{]}}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}\Bigg{)}$ (18)
$\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}({\mathrm{x}}^{(1)})^{2}\Bigg{)}\exp\Bigg{(}\frac{(\sigma r)^{2}}{2}\Big{(}\|{\mathbf{x}}\|^{2}-({\mathrm{x}}^{(1)})^{2}\Big{)}\Bigg{)}$ (19)
$=\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (20)
Equation 17 follows from the standard formula of the moment-generating function of a zero-mean Gaussian. Equation 19 follows because $|{\mathrm{x}}_{1}^{(1)}|$ is SG($\sigma^{2}$) with $\mathbb{E}\big{[}|{\mathrm{x}}^{(1)}|\big{]}=\mu\sigma$ and $\sum_{j=1}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}=\|{\mathbf{x}}\|^{2}$. Since the derivation above holds for any $r\in\mathbb{R}$, we also have:
$\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\big{(}-\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (21)
Next, plugging in equations (20) and (21) into eq. 13, we get:
$\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}\Big{(}(1-p)\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}+p\exp\big{(}-\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Big{)}$ (22)
$=\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}\Bigg{)}\Big{(}1-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Big{)}$ (23)
$\leq\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Bigg{)},$ (24)
where the last step follows because $\exp(z)\geq 1+z$ for all $z\in\mathbb{R}$. Substituting eq. 24 into eq. 12 followed by eq. 11 yields:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}\leq\exp\Bigg{(}n\Bigg{\{}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Bigg{\}}\Bigg{)}.$ (25)
Plugging in eq. 25 into eq. 10, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}\leq\exp\Bigg{(}n\Bigg{\{}\underbrace{\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}}_{:=\nu}\Bigg{\}}\Bigg{)},$ (26)
for any $r<0$. Let $\nu:=\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}$. We will pick $r=-c\Big{(}\frac{\mu{\mathrm{x}}^{(1)}}{\sigma\|{\mathbf{x}}\|^{2}}\Big{)}$ for some $c\in(0,1)$. Putting this into the expression of $\nu$, we get:
$\nu=\mu^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\Bigg{(}\exp\Bigg{(}2c\mu^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}\Bigg{)}-1\Bigg{)}$ (27)
For brevity, let $u=\mu^{2}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Big{)}^{2}$. Note that $u\leq\frac{1}{4}$, as $\mu\leq\frac{1}{2}$ and ${\mathrm{x}}^{(1)}\leq\|{\mathbf{x}}\|$. Rewriting the above equation in terms of $u$, we get:
$\nu=u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}\exp(2cu)-1\big{)}$ (28)
Note that $2cu\leq\frac{1}{2}$ as $c\leq 1$ and $u\leq\frac{1}{4}$. Therefore, applying Lemma 1 above, we get:
$\nu\leq u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}2cu+4c^{2}u^{2}\big{)}\leq\underbrace{u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}2cu+c^{2}u\big{)}}_{\text{(A)}},$ (29)
where the last step is obtained by using the fact that $u\leq\frac{1}{4}$. (A) is minimized by choosing $c=\frac{1-2p}{1+2p}$. With this choice of $c$, we get:
$\nu\leq-\frac{u}{2}\Bigg{(}\frac{(1-2p)^{2}}{1+2p}\Bigg{)}.$ (30)
Further using the fact $1+2p\leq 2$ and plugging in the value of $u=\mu^{2}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Big{)}^{2}$ above, we get:
$\nu\leq-\frac{\mu^{2}}{4}(1-2p)^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}.$ (31)
Plugging this into eq. 26 and recalling that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (32)
Recalling the definition of $\varepsilon$ (eq. 9) and using the result before that in eq. 8, we obtain the following bound for the case of ${\mathrm{y}}=1$:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (33)
This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case, i.e. ${\mathrm{y}}=-1$, can be derived similarly as above by analyzing with $r>0$ instead. In particular, for the case of ${\mathrm{y}}=-1$, we get:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle>0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (34)
(b) Upper bound $\widehat{\alpha}({\mathbf{x}})$. From eq. 8 and eq. 9, recall that:
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle=\frac{\varepsilon}{n\sigma^{2}},\text{ where }\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle.$ (35)
Again, we will prove the result for the case of ${\mathrm{y}}=1$ (i.e., ${\mathrm{x}}^{(1)}>0$); the result for the other case can be derived similarly.
Note that:
$\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\mathrm{x}_{k}^{(1)}\mathrm{x}^{(1)}+\underbrace{\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\sum_{l=2}^{d}\mathrm{x}_{k}^{(l)}\mathrm{x}^{(l)}}_{\tilde{\varepsilon}}.$ (36)
Conditioned on $\widehat{\mathrm{y}}_{k}$ and $\mathrm{x}_{k}^{(1)}$, $\tilde{\varepsilon}\sim\mathcal{N}\big{(}0,n\sigma^{2}\sum_{l=2}^{d}(\mathrm{x}^{(l)})^{2}\big{)}$ using the fact that $\mathrm{x}^{(1)},\mathrm{x}^{(2)},\ldots,\mathrm{x}^{(d)}$ are all independent of each other. But $\sum_{l=2}^{d}(\mathrm{x}^{(l)})^{2}=\|{\mathbf{x}}\|^{2}-(\mathrm{x}^{(1)})^{2}=\|{\mathbf{x}}\|^{2}\Big{(}1-\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}=\|{\mathbf{x}}\|^{2}\textup{sin}^{2}\big{(}\beta({\mathbf{x}})\big{)}$. So in other words, conditioned on $\widehat{\mathrm{y}}_{k}$ and $\mathrm{x}_{k}^{(1)}$, $\tilde{\varepsilon}\sim\mathcal{N}\big{(}0,n\sigma^{2}\|{\mathbf{x}}\|^{2}\textup{sin}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}$. Also, recalling that ${\mathrm{y}}_{k}=\text{sign}(\mathrm{x}_{k}^{(1)})$ and ${\mathrm{x}}^{(1)}>0$, we have that $\widehat{\mathrm{y}}_{k}\mathrm{x}_{k}^{(1)}\mathrm{x}^{(1)}=\widehat{\mathrm{y}}_{k}{\mathrm{y}}_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}$. Let $\delta_{k}:=\widehat{\mathrm{y}}_{k}{\mathrm{y}}_{k}$. Using eq. 1, we have:
$\delta_{k}=\begin{cases}1\quad\text{ w.p. }1-p,\\ -1\quad\text{ o/w.}\end{cases}$ (37)
Moreover, $\delta_{k}$ is independent of $\mathrm{x}_{k}^{(1)}$. Using of all this, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $=\mathbb{P}\Big{(}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}+\varepsilon\leq 0\Big{)}$ (38)
$=\mathbb{E}_{\big{\{}\delta_{k},\mathrm{x}_{k}^{(1)}\big{\}}_{k=1}^{n}}\Bigg{[}\Phi\Bigg{(}-\frac{\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)}\Bigg{]}$ (39)
$\geq\mathbb{E}_{\big{\{}\delta_{k},\mathrm{x}_{k}^{(1)}\big{\}}_{k=1}^{n}}\Bigg{[}\Phi\Bigg{(}-\frac{\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\times\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)}\Bigg{]}.$ (40)
Using the convexity of $\Phi(.)$ on $\mathbb{R}_{\leq 0}$, we can further lower bound eq. 40 by:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $\geq\Phi\Bigg{(}-\frac{\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\times\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)},$ (41)
using Jensen’s inequality. Using the result of Lemma 2 and the fact that $\big{|}\mathrm{x}^{(1)}\big{|}=\|{\mathbf{x}}\|\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}$ above, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $\geq\Phi\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Big{)}.$ (42)
Recalling eq. 35 and using the fact that $\Phi(-z)=\Phi^{\text{c}}(z)$ for all $z\in\mathbb{R}$ (recall that $\Phi^{\text{c}}(.)$ is the complementary CDF of a standard normal variable), we obtain the following bound for the case of ${\mathrm{y}}=1$:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\geq\Phi^{\text{c}}\Big{(}{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Big{)}.$ (43)
Using Formula 7.1.13 from Abramowitz and Stegun, (1968), we have for any $z>0$:
$\Phi^{\text{c}}(z)>\sqrt{\frac{2}{\pi}}\Bigg{(}\frac{e^{-\frac{z^{2}}{2}}}{z+\sqrt{z^{2}+4}}\Bigg{)}.$
Further, using the fact that $z+\sqrt{z^{2}+4}<2(z+1)$ for $z>0$ above, we get:
$\Phi^{\text{c}}(z)>\frac{1}{\sqrt{2\pi}}\Bigg{(}\frac{e^{-\frac{z^{2}}{2}}}{z+1}\Bigg{)}.$ (44)
Now if $z\leq 2$, $z+1\leq 3$. Using this in eq. 44, we get:
$\Phi^{\text{c}}(z)>\frac{1}{3\sqrt{2\pi}}\exp\Big{(}-\frac{z^{2}}{2}\Big{)}.$ (45)
If $z>2$, $z+1<e^{z}<e^{z^{2}/2}$. Using this in eq. 44, we get:
$\Phi^{\text{c}}(z)>\frac{1}{\sqrt{2\pi}}\exp(-{z^{2}}).$ (46)
Combining eq. 45 and eq. 46, we get the following simpler bound:
$\Phi^{\text{c}}(z)>\frac{1}{3\sqrt{2\pi}}\exp(-{z^{2}}).$ (47)
Using this in eq. 43, we get:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\geq\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}.$ (48)
This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case (i.e., ${\mathrm{y}}=-1$) can be derived similarly as above; in particular, here we get:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle>0\Big{)}\geq\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}.$ (49)
This finishes the proof. ∎
Lemma 1.
For $t\leq\frac{1}{2}$, it holds that:
$e^{t}-1\leq t+t^{2}.$
Proof.
Using the Taylor expansion of $e^{t}$, we get:
$e^{t}-1$ $=t+\frac{t^{2}}{2!}+\frac{t^{3}}{3!}+\frac{t^{4}}{4!}\ldots$ (50)
$\leq t+\frac{t^{2}}{2}\big{(}1+t+t^{2}+\ldots\big{)}$ (51)
$=t+\frac{t^{2}}{2(1-t)}.$ (52)
Since $t\leq\frac{1}{2}$, $2(1-t)\geq 1$. Using this above gives us the desired result. ∎
Lemma 2.
For $p\leq\frac{1}{2}$, we have:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\sqrt{n}\sigma+n(1-2p)\mu\sigma.$
Proof.
First, note that:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\sqrt{\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}}.$ (53)
Now using independence, we have:
$\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}\leq\sum_{k}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}^{2}\Big{]}+\sum_{k\neq l}\mathbb{E}\big{[}\delta_{k}\big{]}\mathbb{E}\big{[}\delta_{l}\big{]}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{]}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{l}^{(1)}\big{|}\Big{]}.$ (54)
As per our setting, $\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}^{2}\Big{]}=\sigma^{2}$, $\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{]}=\mu\sigma$ and $\mathbb{E}\big{[}\delta_{k}\big{]}=1-2p$. Using all of this in eq. 54, we get:
$\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}\leq n\sigma^{2}+n(n-1)(1-2p)^{2}\mu^{2}\sigma^{2}\leq n\sigma^{2}+n^{2}(1-2p)^{2}\mu^{2}\sigma^{2}.$ (55)
Using this in eq. 53, we get:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\Big{(}n\sigma^{2}+n^{2}(1-2p)^{2}\mu^{2}\sigma^{2}\Big{)}^{1/2}\leq\sqrt{n}\sigma+n(1-2p)\mu\sigma.$ (56)
This finishes the proof. ∎
Appendix C Proof of Theorem 2
Proof.
Just like Theorem 1, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ without loss of generality. In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$, the ground truth label of ${\mathbf{x}}$ is $\text{sign}({\mathrm{x}}^{(1)})$, $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ and $\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\sqrt{\sum_{j=2}^{d}({\mathrm{x}}^{(j)})^{2}}}$. Also note that:
$\text{acc}(\widehat{\bm{\theta}}_{0})=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Big{]}.$
(a) Lower bound. For the above case, we have:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (57)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (58)
$\geq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Big{[}\alpha({\mathbf{x}})\Big{]}=1-\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]},$ (59)
where eq. 59 follows from Theorem 1.121212To be explicit, in the series of equations above, we can go from eq. 58 to eq. 59 because $\mathcal{X}_{L}$ has zero measure. For brevity, let $t:=\frac{n\mu^{2}}{4}(1-2p)^{2}$. Also, we will omit the subscript ${\mathbf{x}}\sim\mathcal{D}$ henceforth for conciseness. Now, for any $\nu\in(0,1)$:
$\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{]}=\underbrace{\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{|}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{]}}_{\leq 1}\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}\\ +\underbrace{\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{|}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}>\nu\Big{]}}_{\leq\exp(-t\nu^{2})}\underbrace{\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}>\nu\Big{)}}_{\leq 1}.$ (60)
Thus,
$\mathbb{E}\Big{[}\exp\Big{(}-t\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}\Big{]}\leq\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}+\exp(-t\nu^{2}).$ (61)
From Lemma 3, we have for $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$:
$\mathbb{P}\Big{(}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$ (62)
Using this in eq. 61 and plugging in the value of $t$, we get:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\nu^{2}\Bigg{)},$ (63)
for any $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. Since $\gamma<\frac{1}{2}$, we can choose $\nu=\frac{\gamma}{\sqrt{3d}}$. Plugging this choice into eq. 63 yields:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}\Bigg{]}$ $\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{12d}\Bigg{)}.$ (64)
Plugging this into eq. 59 gives us:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $\geq 1-\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}-\exp\Bigg{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{12d}\Bigg{)}$ (65)
$=1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (66)
(b) Upper bound. We have:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (67)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (68)
$\leq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Big{[}\widehat{\alpha}({\mathbf{x}})\Big{]}=1-\frac{1}{3\sqrt{2\pi}}\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}\Bigg{]},$ (69)
where eq. 69 follows from Theorem 1.131313To be explicit, in the series of equations above, we can go from eq. 68 to eq. 69 because $\mathcal{X}_{L}$ has zero measure. For brevity, let $t:=(1+\sqrt{n}\mu(1-2p))^{2}$. Also, we will omit the subscript ${\mathbf{x}}\sim\mathcal{D}$ henceforth for conciseness. Now, for any $\nu>0$:
$\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{]}$ $\geq\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{|}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{]}\mathbb{P}\Big{(}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}$ (70)
$\geq\exp\big{(}-{t\nu^{2}}\big{)}\mathbb{P}\Big{(}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\leq\nu\Big{)}.$ (71)
Let us choose $\nu=\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$. Then, using the result of Lemma 4, we have:
$\mathbb{E}\Big{[}\exp\big{(}-t\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}\Big{]}$ $\geq\frac{1}{2}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}\exp\Big{(}-\frac{8\mu^{2}t}{d-1}\Big{)}.$ (72)
Using this in eq. 69 after plugging in the value of $t=(1+\sqrt{n}\mu(1-2p))^{2}\leq 2(1+n\mu^{2}(1-2p)^{2})$, we get:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $\leq 1-\frac{1}{6\sqrt{2\pi}}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}\exp\Bigg{(}-\frac{16\mu^{2}(1+n\mu^{2}(1-2p)^{2})}{d-1}\Bigg{)}$ (73)
$\leq 1-\Theta(1)\exp\Bigg{(}-\frac{n\mu^{4}(1-2p)^{2}}{\Theta(d)}\Bigg{)}.$ (74)
This completes the proof. ∎
Lemma 3.
In the setting of the proof of Theorem 2, we have for any $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$
Proof.
We have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}$ $=\mathbb{P}\Bigg{(}\frac{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}}{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}+\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}}\geq\nu^{2}\Bigg{)}$ (75)
$=\mathbb{P}\Bigg{(}\sum_{j=2}^{d}\Big{(}\frac{{\mathrm{x}}^{(j)}}{\sigma}\Big{)}^{2}\leq\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\sigma}\Big{)}^{2}\Bigg{)}.$ (76)
Let ${\mathrm{z}}:=\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}/\sigma\big{)}^{2}$ and $\overline{\mathrm{x}}^{(1)}:={\mathrm{x}}^{(1)}/\sigma$. Note that ${\mathrm{z}}$ is a chi-squared random variable with $(d-1)$ degrees of freedom and $\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\geq\gamma$ per our setting. Thus, continuing from eq. 76, we have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}$ $\geq\mathbb{P}\Bigg{(}{\mathrm{z}}<\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\gamma^{2}\Bigg{)}.$ (77)
A simple application of Theorem 1 of Ghosh, (2021) yields:
$\mathbb{P}\big{(}{\mathrm{z}}\geq 3(d-1)\big{)}\leq\exp\Bigg{(}-\frac{(d-1)}{2}(2-\log 3)\Bigg{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}.$ (78)
Thus, for any $w\geq 3(d-1)$, $\mathbb{P}({\mathrm{z}}<w)\geq 1-\exp\Big{(}-\frac{9(d-1)}{20}\Big{)}$. Using this in eq. 77 yields:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\geq\nu\Big{)}\geq 1-\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)},$ (79)
when $\Big{(}\frac{1-\nu^{2}}{\nu^{2}}\Big{)}\gamma^{2}\geq 3(d-1)$ or $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. So:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cos}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)},$ (80)
for $\nu\leq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}$. This finishes the proof. ∎
Lemma 4.
In the setting of the proof of Theorem 2, we have for any $\nu\geq\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}\geq\frac{1}{2}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}.$
Proof.
We have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $=\mathbb{P}\Bigg{(}\frac{\big{(}{\mathrm{x}}^{(1)}\big{)}^{2}}{\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}\big{)}^{2}}\leq\nu^{2}\Bigg{)}$ (81)
$=\mathbb{P}\Bigg{(}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\sigma}\Big{)}^{2}\leq\nu^{2}\sum_{j=2}^{d}\Big{(}\frac{{\mathrm{x}}^{(j)}}{\sigma}\Big{)}^{2}\Bigg{)}.$ (82)
Let ${\mathrm{z}}:=\sum_{j=2}^{d}\big{(}{\mathrm{x}}^{(j)}/\sigma\big{)}^{2}$ and $\overline{\mathrm{x}}^{(1)}:={\mathrm{x}}^{(1)}/\sigma$. Note that ${\mathrm{z}}$ is a chi-squared random variable with $(d-1)$ degrees of freedom. Thus, continuing from eq. 82, we have:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $\geq\mathbb{P}\Big{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}{\mathrm{z}}\Big{)}$ (83)
$\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}{\mathrm{z}}\text{ }\big{|}\text{ }{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}$ (84)
$\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}\Big{(}\frac{d-1}{2}\Big{)}\Bigg{)}\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}.$ (85)
From Theorem 2 of Ghosh, (2021), we have that:
$\mathbb{P}\Bigg{(}{\mathrm{z}}>\frac{d-1}{2}\Bigg{)}\geq 1-\exp\Bigg{(}-\frac{d-1}{16}\Bigg{)}.$ (86)
Plugging this into eq. 85, we get:
$\mathbb{P}\Big{(}{\mathbf{x}}:|\textup{cot}(\beta({\mathbf{x}}))|\leq\nu\Big{)}$ $\geq\mathbb{P}\Bigg{(}\big{(}\overline{\mathrm{x}}^{(1)}\big{)}^{2}\leq\nu^{2}\Big{(}\frac{d-1}{2}\Big{)}\Bigg{)}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}$ (87)
$=\mathbb{P}\Bigg{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq\nu\sqrt{\frac{d-1}{2}}\Bigg{)}\Bigg{(}1-\exp\Big{(}-\frac{d-1}{16}\Big{)}\Bigg{)}.$ (88)
As per our setting, we have that $\mathbb{E}[\big{|}\overline{\mathrm{x}}^{(1)}\big{|}]=\mu$. Note that for $\nu\geq\frac{2\sqrt{2}\mu}{\sqrt{d-1}}$, we have that:
$\mathbb{P}\Bigg{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq\nu\sqrt{\frac{d-1}{2}}\Bigg{)}\geq\mathbb{P}\Big{(}\big{|}\overline{\mathrm{x}}^{(1)}\big{|}\leq 2\mu\Big{)}\geq\frac{1}{2},$ (89)
where the last step above follows from Markov’s inequality. Plugging in eq. 89 into eq. 88 gives us the desired result. ∎
Appendix D Proof of Theorem 3
Proof.
[…]
We first follow a standard argument to “reduce” the classification problem to a multi-way hypothesis testing problem. Let $\rho\in(0,1)$ be an arbitrary but fixed value which we can choose. We define a $\rho$-packing of $S^{d-1}$ as a set $\Theta=\{\bm{\theta}_{1},\dotsc,\bm{\theta}_{M}\}\subset S^{d-1}$ such that $\langle\bm{\theta}_{l},\bm{\theta}_{k}\rangle\leq\rho$ for $l\neq k$. Also define the $\rho$-packing number of $S^{d-1}$ as
$M(\rho,S^{d-1}):=\sup\{M\in\mathbb{N}:\text{there exists a $\rho$-packing $\Theta$ of $S^{d-1}$ with size $M$}\}\,.$ (90)
For convenience, we define the misclassification error of a classifier $\widehat{\bm{\theta}}\in S^{d-1}$ as $\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})=1-\textup{acc}(\widehat{\bm{\theta}};\bm{\theta}^{*})$. By our assumption,
$\delta\geq\sup_{\bm{\theta}^{\ast}\in S^{d-1}}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})\geq\sup_{\bm{\theta}^{\ast}\in\Theta}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})\,.$ (91)
In order to further lower bound the right hand side, we let $I$ be a random variable uniformly distributed on the hypothesis set $\{1,2,\dotsc,M\}$ and consider the case of $\bm{\theta}^{\ast}=\bm{\theta}_{I}$. We also define $\widehat{I}$ as the index of the element in $\Theta$ with maximum inner product with $\widehat{\bm{\theta}}$ (it does not matter how we break ties). We then have
$\sup_{\bm{\theta}^{\ast}\in\Theta}\textup{err}(\widehat{\bm{\theta}};\bm{\theta}^{*})$ $\geq\max_{i\in[M]}\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$
$\geq\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$
$\geq\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}(\widehat{I}\neq i|I=i)$ (92)
$=\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\mathbb{P}(\widehat{I}\neq I)\,,$ (93)
with $\Phi$ denoting the CDF of a standard normal variable. Equation 92 above follows from the lemma below.
Lemma 5.
For any $i\in[M]$, we have
$\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)\geq\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\;\mathbb{P}(\widehat{I}\neq i|I=i)\,.$
Combining (91) and (93), we obtain that
$\delta_{0}:=\frac{\delta}{\Phi\Big{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Big{)}}\geq\mathbb{P}(\widehat{I}\neq I)\,.$ (94)
Next recall the set $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$, and let $\mathbf{X}=[{\mathbf{x}}_{1}^{T},\dotsc,{\mathbf{x}}_{n}^{T}]^{T}\in\mathbb{R}^{n\times d}$ and $\widehat{\mathbf{y}}=[\widehat{\mathrm{y}}_{1},\dotsc,\widehat{\mathrm{y}}_{n}]^{T}$. By an application of Fano’s inequality, with conditioning on $\mathbf{X}$ (see e.g. (Scarlett and Cevher,, 2019, Section 2.3)) we have
$\mathcal{I}(I;\widehat{I}|\mathbf{X})\geq\left(1-\delta_{0}\right)\log(M(\rho,S^{d-1}))-\log 2\,,$ (95)
where $\mathcal{I}(I;\widehat{I}|\mathbf{X})$ represents the conditional mutual information between $I$ and $\widehat{I}$. Using the fact that $I\to\widehat{\mathbf{y}}\to\widehat{I}$ forms a Markov chain conditioned on $\mathbf{X}$, and by an application of the data processing inequality we have:
$\mathcal{I}(I;\widehat{I}|\mathbf{X})\leq\mathcal{I}(I;\widehat{\mathbf{y}}|\mathbf{X})$ (96)
We will now upper bound $\mathcal{I}(I;\widehat{\mathbf{y}}|\mathbf{X})$. Let $\mathrm{w}_{j}:=\frac{1}{2}\big{(}\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)+1\big{)}$ be the 0-1 version of the actual label of $\mathbf{x}_{j}$, viz., $\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)$. As per our setting, we have $\widehat{\mathrm{y}}_{j}=2\big{(}\mathrm{w}_{j}\oplus z_{j}\big{)}-1$ where $z_{j}\sim\text{Bernoulli}(p)$ and $\oplus$ denotes modulo-2 addition. Since the noise variables $z_{j}$ are independent and $\widehat{\mathrm{y}}_{j}$ depends on $(I,\mathbf{X})$ only through $\mathrm{w}_{j}=\frac{1}{2}\big{(}\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)+1\big{)}$, by using the tensorization property of the mutual information (see e.g. (Scarlett and Cevher,, 2019, Lemma 2, part (iii))), we have
$\mathcal{I}(I;\mathbf{y}|\mathbf{X})\leq\sum_{j=1}^{n}\mathcal{I}\big{(}\mathrm{w}_{j};\widehat{\mathrm{y}}_{j}\big{)}\leq n(\log 2-H_{2}(p))\,,$ (97)
where the second inequality follows since $\widehat{\mathrm{y}}_{j}$ is generated by passing $\mathrm{w}_{j}$ through a binary symmetric channel, which has capacity $\log 2-H_{2}(p)$ with $H_{2}(p):=-p\log p-(1-p)\log(1-p)$ denoting the binary entropy function.
We next use the lemma below to further upper bound the right-hand side of (97).
Lemma 6.
For a discrete probability distribution, consider the entropy function given by
$H(p_{1},\dotsc,p_{k})=\sum_{i=1}^{k}p_{k}\log(1/p_{k})\,.$
We have the following bound:
$H(p_{1},\dotsc,p_{k})\geq\log k-k\sum_{i=1}^{k}(p_{i}-1/k)^{2}\,.$
Using Lemma 6 with $k=2$ we obtain $\log 2-H_{2}(p)\leq 4(p-1/2)^{2}=(1-2p)^{2}$, which along with (97), (96) and (95) gives
$\frac{n(1-2p)^{2}+\log 2}{1-\delta_{0}}\geq\log(M(\rho,S^{d-1}))\,.$ (98)
In our next lemma, we lower bound $M(\rho,S^{d-1})$.
Lemma 7.
Recall the definition of $\rho$-packing number of $S^{d-1}$ given by (90). We have the following bound:
$M(\rho,S^{d-1})\geq\exp\Big{(}\frac{d\rho^{2}}{2}\Big{)}\,.$
Using Lemma 7 along with (98), we obtain the following lower bound on the sample complexity:
$n\geq\frac{\frac{\rho^{2}}{2}(1-\delta_{0})d-\log 2}{(1-2p)^{2}},\text{ with }\delta_{0}=\frac{\delta}{\Phi\Big{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Big{)}}.$
Let us pick $\rho=\frac{(4/5)^{2}-\mu^{2}}{(4/5)^{2}+\mu^{2}}$; it can be shown that $\rho>\frac{2}{5}$, because $\mu\leq\frac{1}{2}$ as per our setting. In that case, using the fact that $\Phi(-\frac{4}{5})>\frac{1}{5}$, we have that $\delta_{0}<5\delta$. Using all of this gives us the desired result for $\delta<\frac{1}{5}$.
We now prove Lemmas 5, 6 and 7.
Proof of Lemma 5. Let us consider the event $\widehat{I}\neq i$, given that $I=i$. We note that if $\widehat{I}\neq I$, then $\langle\widehat{\bm{\theta}},\bm{\theta}_{I}\rangle\leq\sqrt{\frac{1+\rho}{2}}$ or equivalently, the angle between $\widehat{\bm{\theta}}$ and $\bm{\theta}_{I}$ is $\geq b:=\cos^{-1}\Big{(}\sqrt{\frac{1+\rho}{2}}\Big{)}$. Otherwise, by definition of $\widehat{I}$ we have $\langle\widehat{\bm{\theta}},\bm{\theta}_{\widehat{I}}\rangle\geq\langle\widehat{\bm{\theta}},\bm{\theta}_{I}\rangle>\sqrt{\frac{1+\rho}{2}}$, and therefore the angle between $\bm{\theta}_{\widehat{I}}$ and $\bm{\theta}_{I}$ is $<2b$. Noting that $\cos(2b)=2\cos^{2}(b)-1=\rho$, we would then have $\langle\bm{\theta}_{\widehat{I}},\bm{\theta}_{I}\rangle>\rho$, which is a contradiction since $\Theta$ forms a $\rho$-packing.
Next, under our data model and conditioned on the value of $I$, $\mathbb{E}\big{[}|\langle\mathbf{x},\bm{\theta}_{I}\rangle|\big{]}=\mu$ and on the space orthogonal to $\bm{\theta}_{I}$, $\mathbf{x}$ follows $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d-1})$ distribution. By rotating the basis and without loss of generality in lower bounding $\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$, we can assume $\bm{\theta}_{I}=\bm{\theta}_{i}=\mathbf{e}_{1}$ and $\widehat{\bm{\theta}}$ is an arbitrary vector in $S^{d-1}$ such that $\theta:=\langle\widehat{\bm{\theta}},\bm{e}_{1}\rangle\leq\sqrt{\frac{1+\rho}{2}}$. We proceed by writing
$\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)=\mathbb{P}\left(\mathrm{sign}({\mathrm{x}}^{(1)})\Big{(}{\mathrm{x}}^{(1)}\theta+\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\Big{)}\leq 0\right)\,,$
where for a vector ${\mathbf{v}}\in\mathbb{R}^{d}$, we use the notation ${\mathbf{v}}^{(-1)}\in\mathbb{R}^{d-1}$ to refer to the vector obtained by dropping its first coordinate. Note that $\mathbf{x}$ is a test data point, independent of the training data $\mathcal{T}$ and so it is independent of $\bm{\widehat{\theta}}$. Therefore, we have $\mathrm{sign}({\mathrm{x}}^{(1)})\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\sim\mathcal{N}(0,\sigma^{2}\|\widehat{\bm{\theta}}^{(-1)}\|^{2})$. By invoking the condition $\|\widehat{\bm{\theta}}\|=1$, this can be written as $\mathrm{sign}({\mathrm{x}}^{(1)})\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle=\sqrt{1-\theta^{2}}\,Z$, where $Z$ is a standard normal variable. Hence, we have:
$\mathbb{P}\left(\mathrm{sign}({\mathrm{x}}^{(1)})\Big{(}{\mathrm{x}}^{(1)}\theta+\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\Big{)}\leq 0\right)$
$=\mathbb{P}\left(|{\mathrm{x}}^{(1)}|\theta+\sqrt{1-\theta^{2}}\,Z\leq 0\right)$
$=\mathbb{P}\left(Z\leq-\frac{|{\mathrm{x}}^{(1)}|\theta}{\sqrt{1-\theta^{2}}}\right)$
$=1-\mathbb{E}\Bigg{[}\Phi\Big{(}\frac{\theta}{\sqrt{1-\theta^{2}}}|{\mathrm{x}}^{(1)}|\Big{)}\Bigg{]}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}1-\mathbb{E}\Bigg{[}\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}|{\mathrm{x}}^{(1)}|\Bigg{)}\Bigg{]}$
$\stackrel{{\scriptstyle(b)}}{{\geq}}1-\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}\mathbb{E}\big{[}|{\mathrm{x}}^{(1)}|\big{]}\Bigg{)}$
$=1-\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}$
$=\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\,,$
where $(a)$ follows from the fact that $\theta\leq\sqrt{\frac{1+\rho}{2}}$ and $(b)$ holds due to Jensen’s inequality and concavity of $\Phi(.)$ on the positive values. Combining the above bound with the event $\widehat{I}\neq i$ given that $I=i$ gives us the desired result.
Proof of Lemma 6. Define $q_{i}=p_{i}-1/k$. Note that $q_{i}$ can be negative, and we have $\sum_{i=1}^{k}q_{i}=0$. We write
$H(p_{1},\dotsc,p_{k})$ $=-\sum_{i=1}^{k}p_{i}\log p_{i}$
$=-\sum_{i=1}^{k}(1/k+q_{i})\log(1/k+q_{i})$
$=-\sum_{i=1}^{k}(1/k+q_{i})[\log(1/k)+\log(1+kq_{i})]$
$\geq\log k-\sum_{i=1}^{k}(1/k+q_{i})kq_{i}$ (99)
$=\log k-k\sum_{i=1}^{k}q_{i}^{2}$
$=\log k-k\sum_{i=1}^{k}(p_{i}-1/k)^{2}\,.$
Note that in eq. 99 we used the fact that $1+kq_{i}\geq 0$ and $\log x\leq x-1$ for all $x\geq 0$.
This completes the proof of the lemma.
Proof of Lemma 7. Define a $\rho$-cover of $S^{d-1}$ as a set of $\mathcal{V}:=\{\mathbf{v}_{1},\dotsc,\mathbf{v}_{N}\}$ such that for any $\bm{\theta}\in S^{d-1}$, there exists some $\mathbf{v}_{i}$ such that $\langle\bm{\theta},\mathbf{v}_{i}\rangle\geq\rho$. The $\rho$-covering number of $S^{d-1}$ is
$N(\rho,S^{d-1}):=\inf\{N\in\mathbb{N}:\text{there exists a $\rho$-cover $\mathcal{V}$ of $S^{d-1}$ with size $N$}\}\,.$
By a simple argument we have $M(\rho,S^{d-1})\geq N(\rho,S^{d-1})$. Concretely, we construct a $\rho$-packing greedily by adding an element at each step which has inner product at most $\rho$ with all the previously selected elements, until it is no longer possible. This means that any point on $S^{d-1}$ has inner product larger than $\rho$ by some of the elements in the constructed set (otherwise it contradicts its maximality). Hence, we have a set that is both a $\rho$-cover and a $\rho$-packing of $S^{d-1}$, and by definition it results in $M(\rho,S^{d-1})\geq N(\rho,S^{d-1})$.
We next lower bound $N(\rho,S^{d-1})$ via a volumetric argument. Let $\mathcal{V}:=\{\mathbf{v}_{1},\dotsc,\mathbf{v}_{N}\}$ be a $\rho$-cover of $S^{d-1}$. For each element $\mathbf{v}_{i}\in\mathcal{V}$ we consider the cone around it with apex angle $\textup{cos}^{-1}(\rho)$. Its intersection with $S^{d-1}$ defines a spherical cap which we denote by $\mathcal{C}(\mathbf{v}_{i},\rho)$. Since $\mathcal{V}$ forms a $\rho$-cover of $S^{d-1}$, we have
$\textup{Vol}(S^{d-1})\leq\textup{Vol}(\cup_{i=1}^{N}\mathcal{C}(\mathbf{v}_{i},\rho))\leq\sum_{i=1}^{N}\textup{Vol}(\mathcal{C}(\mathbf{v}_{i},\rho))\,.$
We next use Lemma 2.2 from Ball et al., (1997) by which we have $\frac{\mathcal{C}(\mathbf{v}_{i},\rho)}{\textup{Vol}(S^{d-1})}\leq e^{-d\rho^{2}/2}$. Using this above, we get
$1\leq Ne^{-d\rho^{2}/2}\,$
for any $\rho$-cover $\mathcal{V}$. Thus, we have $N(\rho,S^{d-1})\geq\exp(\frac{d\rho^{2}}{2})$, which completes the proof of the lemma.
Appendix E Proof of Theorem 4
Proof.
Recall the problem setting at the beginning of Section 4. Without loss of generality, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ (just like the proof of Theorem 1). In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$.
Let ${\mathrm{y}}=\text{sign}({\mathrm{x}}^{(1)})$ be the ground truth label of ${\mathbf{x}}$. Note that proving Theorem 4 is equivalent to proving the following statement with $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}$. For a sample ${\mathbf{x}}\notin\mathcal{X}_{L}\cup\mathcal{X}_{U}$ such that $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$, the following holds with a probability of at least $1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$:
$\begin{cases}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\geq 0\text{ when }{\mathrm{y}}=1\text{ i.e., }{\mathrm{x}}^{(1)}\geq 0\\ \langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle<0\text{ when }{\mathrm{y}}=-1\text{ i.e., }{\mathrm{x}}^{(1)}<0.\end{cases}$ (100)
Also note that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ in our choice of basis for the proof.
Just like the proof of Theorem 1, we will prove the above for the case of ${\mathrm{y}}=1$ or equivalently $\mathrm{x}^{(1)}\geq 0$, and the result for the other case (i.e., ${\mathrm{y}}=-1$ or equivalently $\mathrm{x}^{(1)}<0$) can be derived similarly.
Recall that:
$\widehat{\bm{\theta}}_{1}:=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widetilde{\mathrm{y}}_{j}^{\prime}{\mathbf{x}}_{j}^{\prime}\Bigg{)}.$ (101)
Note that:
$\widetilde{\mathrm{y}}_{j}^{\prime}=\begin{cases}\mathrm{y}_{j}^{\prime}\text{w.p. }\geq\alpha({\mathbf{x}}_{j}^{\prime})\\ -\mathrm{y}_{j}^{\prime}\text{ w.p. }\leq 1-\alpha({\mathbf{x}}_{j}^{\prime}),\end{cases}$ (102)
where $\alpha({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ is as defined in Theorem 1.
We have:
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle=\frac{1}{\sigma^{2}}\Bigg{\{}\frac{1}{n}\sum_{k=1}^{n}\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle\Bigg{\}}.$ (103)
Since $\frac{1}{n\sigma^{2}}$ is a constant independent of $\mathcal{T}$ and $\mathcal{T}_{2}$, we will analyze
$\varepsilon^{\prime}=\sum_{k=1}^{n}\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle.$ (104)
Using the Chernoff bound, we have for any $r<0$:
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}$ $=\mathbb{P}\big{(}\exp\big{(}r\varepsilon^{\prime}\big{)}\geq 1\big{)}\leq\mathbb{E}\Big{[}\exp\big{(}r\varepsilon^{\prime}\big{)}\Big{]}.$ (105)
For $k\in[n]$, let $\varepsilon_{k}^{\prime}:=\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle$; thus, $\varepsilon^{\prime}=\sum_{k\in[n]}\varepsilon_{k}^{\prime}$. For any $r\in\mathbb{R}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon^{\prime}\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r\sum_{k\in[n]}\varepsilon_{k}^{\prime}\big{)}\Big{]}=\mathbb{E}\Big{[}\prod_{k\in[n]}\exp\big{(}r\varepsilon_{k}^{\prime}\big{)}\Big{]}=\Big{(}\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}^{\prime}\big{)}\Big{]}\Big{)}^{n},$ (106)
where the last step follows because of the i.i.d. nature of the data. Now:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}^{\prime}\big{)}\Big{]}$ $=\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}.$ (107)
Taking expectation w.r.t. the randomness in $\widetilde{\mathrm{y}}_{1}^{\prime}$ first, we get:
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Bigg{[}\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\big{|}\widetilde{\mathrm{y}}_{1}^{\prime}={\mathrm{y}}_{1}^{\prime}\Big{]}\underbrace{\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}={\mathrm{y}}_{1}^{\prime}\Big{)}}_{=1-\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}}\\ +\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\big{|}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{]}\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{)}\Bigg{]}.$ (108)
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}+\underbrace{\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{)}\Bigg{]}}_{:=\text{(I)}}.$ (109)
We already evaluated $\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ in the proof of Theorem 1; see eq. 20. It was:
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (110)
[…]
From eq. 102, we have $\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\leq 1-\alpha({\mathbf{x}}_{1}^{\prime})=\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Big{)}$. Thus:
$\mathbb{E}\Bigg{[}\Big{(}\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\Big{)}^{2}\Bigg{]}\leq\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{2}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Bigg{)}\Bigg{]}.$ (112)
Similar to eq. 64 in the proof of Theorem 2, it can be shown that:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{2}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Bigg{)}\Bigg{]}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)},$ (113)
and so:
$\mathbb{E}\Bigg{[}\Big{(}\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\Big{)}^{2}\Bigg{]}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)}.$ (114)
Next,
$\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}^{2}\Bigg{]}$ $=\mathbb{E}\Big{[}\exp\big{(}-2r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}+\exp\big{(}2r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-2\Big{]}.$ (115)
Using eq. 20 and eq. 21 in the proof of Theorem 1 above, we get:
$\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}^{2}\Bigg{]}\leq\\ \Big{(}\underbrace{\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}+\exp\big{(}2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}}_{=2\text{cosh}(2\mu\sigma r{\mathrm{x}}^{(1)})}\Big{)}\exp\Big{(}2{(\sigma r\|{\mathbf{x}}\|)^{2}}\Big{)}-2.$ (116)
For brevity, let $h(r):=2\text{cosh}\big{(}2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\big{(}2{(\sigma r\|{\mathbf{x}}\|)^{2}}\big{)}-2$. Plugging in eq. 114 and eq. 116 into eq. 111 yields:
(I) $\leq\sqrt{h(r)}\sqrt{\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)}}$
$\leq\sqrt{h(r)}\underbrace{\Bigg{(}\exp\Bigg{(}-\frac{9(d-1)}{40}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Bigg{)}\Bigg{)}}_{=p_{2}},$ (117)
where the last step is obtained using the fact that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for any $a,b\geq 0$. Also, let us put $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Big{)}$. Here we consider the case of $n$ and $d$ being large enough so that $p_{2}\leq\frac{1}{2}$. Now plugging in eq. 110 and eq. 117 into eq. 109 gives us:
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}+p_{2}\sqrt{h(r)}.$ (118)
Let us pick $r=r^{*}:=-\frac{\mu{\mathrm{x}}^{(1)}(1-2p_{2})}{2\sigma\|{\mathbf{x}}\|^{2}}$ (recall ${\mathrm{x}}^{(1)}>0$). Note that $r^{*}<0$ and we needed $r$ to be $<0$ for the Chernoff bound in eq. 105 to hold. With this value of $r$, we have:
$\exp\big{(}\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}$ $=\exp\Bigg{(}-\frac{(\mu{\mathrm{x}}^{(1)})^{2}(1-2p_{2})}{2\|{\mathbf{x}}\|^{2}}\Bigg{(}1-\frac{1-2p_{2}}{4}\Bigg{)}\Bigg{)}$
$\leq\exp\Bigg{(}-\frac{3(\mu{\mathrm{x}}^{(1)})^{2}(1-2p_{2})}{8\|{\mathbf{x}}\|^{2}}\Bigg{)}.$ (119)
Let $u:=\frac{(\mu{\mathrm{x}}^{(1)})^{2}}{\|{\mathbf{x}}\|^{2}}$. The above equation in terms of $u$ is:
$\exp\big{(}\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}{\exp\Bigg{(}\frac{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}}\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}.$ (120)
Also, recall that:
$h(r^{*})$ $=2\text{cosh}\big{(}2\mu\sigma r^{*}{\mathrm{x}}^{(1)}\big{)}\exp\big{(}2{(\sigma r^{*}\|{\mathbf{x}}\|)^{2}}\big{)}-2.$
It can be checked that:
$h(r^{*})={2\text{cosh}\big{(}u(1-2p_{2})\big{)}\exp\Big{(}\frac{u}{2}(1-2p_{2})^{2}\Big{)}-2}.$ (121)
Since $\mu\leq\frac{1}{2}$ and ${\mathrm{x}}^{(1)}\leq\|{\mathbf{x}}\|$, $u\leq\frac{1}{4}$ and thus, $u(1-2p_{2})\leq\frac{1}{4}$ and $\frac{u}{2}(1-2p_{2})^{2}\leq\frac{1}{8}$. Using Lemma 8 and Lemma 9 above, we get:
$h(r^{*})$ $\leq 2\Big{(}1+\frac{8}{15}u^{2}(1-2p_{2})^{2}\Big{)}\Big{(}1+\frac{4}{7}u(1-2p_{2})^{2}\Big{)}-2$ (122)
$={8}u(1-2p_{2})^{2}\Bigg{(}\frac{1}{7}+\frac{2}{15}u+\frac{8}{105}u^{2}(1-2p_{2})^{2}\Bigg{)}.$ (123)
Further, using the fact that $u\leq\frac{1}{4}$ and $u^{2}(1-2p_{2})^{2}\leq\frac{1}{16}$ above, we get:
$h(r^{*})\leq{2u}(1-2p_{2})^{2}.$ (124)
Plugging in eq. 120 and eq. 124 into eq. 118, we get for $r=r^{*}$:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}+{p_{2}\sqrt{2u}}(1-2p_{2})$
$=\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}\Bigg{(}1+{p_{2}\sqrt{2u}}(1-2p_{2})\exp\Big{(}\frac{3u}{8}(1-2p_{2})\Big{)}\Bigg{)}.$
Now note that $\exp\Big{(}\frac{3u}{8}(1-2p_{2})\Big{)}\leq e^{\frac{3}{32}}$ as $u(1-2p_{2})\leq\frac{1}{4}$. Using this above together with the fact that $\sqrt{2}\times e^{\frac{3}{32}}\leq 2$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})\Big{)}\Big{(}1+2{p_{2}\sqrt{u}}(1-2p_{2})\Big{)}$
$\leq\exp\Big{(}-\frac{3u}{8}(1-2p_{2})+2{p_{2}\sqrt{u}}(1-2p_{2})\Big{)},$ (125)
where the last step follows by using the fact that $\exp(z)\geq 1+z$ for all $z\in\mathbb{R}$. Plugging in the value of $u=\frac{(\mu{\mathrm{x}}^{(1)})^{2}}{\|{\mathbf{x}}\|^{2}}$ above and recalling that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$, we get:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Bigg{(}-(1-2p_{2})\Bigg{(}\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\Bigg{)}\Bigg{)}.$ (126)
Plugging in eq. 126 into eq. 107 and then applying that into eq. 106 yields:
$\mathbb{E}\Big{[}\exp\big{(}r^{*}\varepsilon^{\prime}\big{)}\Big{]}\leq\exp\Bigg{(}-n(1-2p_{2})\Bigg{(}\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}\Bigg{)}\Bigg{)}.$ (127)
Next, plugging this into eq. 105 gives us:
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}\leq\exp\Bigg{(}-n(1-2p_{2})\Bigg{(}\underbrace{\frac{3\mu^{2}}{8}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}-{2p_{2}\mu}\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}_{:=\text{(II)}}\Bigg{)}\Bigg{)}.$ (128)
[…]
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (129)
Recalling the definition of $\varepsilon^{\prime}$ (eq. 104) and using the result above that in eq. 103, we obtain the following bound for the case of ${\mathrm{y}}=1$:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\leq 0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)},$ (130)
as long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$ with $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Big{)}$. This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case, i.e. ${\mathrm{y}}=-1$, can be derived similarly as above by analyzing with $r>0$ instead. In particular, for the case of ${\mathrm{y}}=-1$, we get:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle>0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)},$ (131)
again as long as $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$. ∎
Lemma 8.
For $t\leq\frac{1}{4}$, it holds that:
$\textup{cosh}(t)\leq 1+\frac{8t^{2}}{15}.$
Proof.
Using the Taylor expansion of $\textup{cosh}(t)$, we get:
$\textup{cosh}(t)$ $=1+\frac{t^{2}}{2!}+\frac{t^{4}}{4!}+\ldots$ (132)
$\leq 1+\frac{t^{2}}{2}\Big{(}1+t^{2}+\ldots\Big{)}$ (133)
$=1+\frac{t^{2}}{2(1-t^{2})}$ (134)
Since $t\leq\frac{1}{4}$, $1-t^{2}\geq\frac{15}{16}$. Using this above gives us the desired result. ∎
Lemma 9.
For $t\leq\frac{1}{8}$, it holds that:
$e^{t}\leq 1+\frac{8t}{7}.$
Proof.
Using the Taylor expansion of $e^{t}$, we get:
$e^{t}$ $=1+t+\frac{t^{2}}{2!}+\frac{t^{3}}{3!}+\frac{t^{4}}{4!}+\ldots$ (135)
$\leq 1+t\big{(}1+t+t^{2}+\ldots\big{)}$ (136)
$=1+\frac{t}{(1-t)}.$ (137)
Since $t\leq\frac{1}{8}$, $(1-t)\geq\frac{7}{8}$. Using this above gives us the desired result. ∎
Appendix F Proof of Theorem 5
Proof.
Just like the proof of Theorem 2, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ without loss of generality. In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$, the ground truth label of ${\mathbf{x}}$ is $\text{sign}({\mathrm{x}}^{(1)})$ and $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$.
For the above case, we have: [^14] To be explicit, in the series of equations below, we can go from eq. 140 to eq. 141 because $\mathcal{X}_{L}\cup\mathcal{X}_{U}$ has zero measure.
$\text{acc}(\widehat{\bm{\theta}}_{1})$ $=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\text{sign}({\mathrm{x}}^{(1)})\Big{)}$ (138)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T},\mathcal{T}_{2}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (139)
$=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\mathbb{P}_{\mathcal{T},\mathcal{T}_{2}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (140)
$\geq\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\alpha_{1}({\mathbf{x}})\Big{|}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{]}\mathbb{P}\Bigg{(}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{)},$ (141)
where $\alpha_{1}({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ is as defined in Theorem 4 and we will ensure $\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\geq\frac{{16p_{2}}}{\mu}$, which can be simplified to:
$p_{2}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}.$
Recall that $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}=\exp\big{(}-\Theta(d)\big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}$. Note that $\exp\big{(}-\Theta(d)\big{)}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}$ when $\gamma\geq\Theta\Big{(}\frac{\sqrt{d}\exp(-\Theta(d))}{\mu}\Big{)}$. Imposing $\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}(1-2p)^{2}}{\Theta(d)}\Big{)}\leq\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}$ gives us:
$n\geq\Theta\Bigg{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\Bigg{)}.$ (142)
[…]
$n\geq\Theta\Bigg{(}\frac{1}{(1-2p)^{2}}\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)}\max\Bigg{(}\log\Big{(}\frac{d}{\mu^{2}\gamma^{2}}\Big{)},\log\frac{1}{p}\Bigg{)}\Bigg{)}.$ (143)
So our final bound for $p_{2}$ is:
$p_{2}\leq\min\Big{(}p,\Theta\Big{(}\frac{\gamma\mu}{\sqrt{d}}\Big{)}\Big{)}.$ (144)
With the above constraint, we get after applying Lemma 3:
$\mathbb{P}\Bigg{(}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{)}\geq 1-\exp\big{(}-\Theta(d)\big{)}.$ (145)
Moreover, with our constraint:
$\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D}}\Bigg{[}\alpha_{1}({\mathbf{x}})\Big{|}|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{\gamma}{\sqrt{\gamma^{2}+{3}(d-1)}}\Bigg{]}\geq 1-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)}.$ (146)
Plugging in eq. 145 and eq. 146 into eq. 141 gives us:
$\text{acc}(\widehat{\bm{\theta}}_{1})\geq 1-\exp\big{(}-\Theta(d)\big{)}-\exp\Bigg{(}-\frac{n{\mu^{2}\gamma^{2}}(1-2p_{2})}{\Theta(d)}\Bigg{)}.$ (147)
This completes the proof. ∎
Appendix G Roadmap of Analysis with Exact Minimizer $\bar{\bm{\theta}}_{0}$
Let
${\mathbf{P}}:=\Bigg{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}-\sigma^{2}{\mathbf{I}}_{d}\Bigg{)}.$ (148)
Then, we can write:
$\bar{\bm{\theta}}_{0}=\Bigg{(}\frac{1}{n}\sum_{k=1}^{n}{\mathbf{x}}_{k}{\mathbf{x}}_{k}^{\top}\Bigg{)}^{-1}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}=\frac{1}{\sigma^{2}}\Bigg{(}{\mathbf{I}}_{d}+\frac{{\mathbf{P}}}{\sigma^{2}}\Bigg{)}^{-1}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)}.$ (149)
We can bound $\|{\mathbf{P}}\|$ with high probability. For instance, using Theorem 13.3 of Rinaldo and Neopane, (2018), we have with a probability of at least $1-\delta$:
$\|{\mathbf{P}}\|\leq C\sigma^{2}\max\Bigg{(}\sqrt{\frac{d+\log(2/\delta)}{n}},\frac{d+\log(2/\delta)}{n}\Bigg{)},$ (150)
for some absolute constant $C>0$. We will focus on the case of $n>\mathcal{O}\big{(}d+\log(2/\delta)\big{)}$; in that case, we have:
$\|{\mathbf{P}}\|\leq C\sigma^{2}\sqrt{\frac{d+\log(2/\delta)}{n}},$ (151)
with a probability of at least $1-\delta$. From eq. 2, recall:
$\widehat{\bm{\theta}}_{0}=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widehat{\mathrm{y}}_{j}{\mathbf{x}}_{j}\Bigg{)},$ (152)
Then using Lemma 10, we get:
$\|\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\|\leq\Bigg{(}\frac{\frac{\|{\mathbf{P}}\|}{\sigma^{2}}}{1-\frac{\|{\mathbf{P}}\|}{\sigma^{2}}}\Bigg{)}\|\widehat{\bm{\theta}}_{0}\|.$ (153)
Next, using eq. 151 above, we get:
$\|\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\|\leq\Bigg{(}\frac{C\sqrt{\frac{d+\log(2/\delta)}{n}}}{1-{C}\sqrt{\frac{d+\log(2/\delta)}{n}}}\Bigg{)}\|\widehat{\bm{\theta}}_{0}\|,$ (154)
with a probability of at least $1-\delta$. Note that:
$\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}\rangle$ $=\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle+\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}-\widehat{\bm{\theta}}_{0}\rangle.$
Using the Cauchy-Schwarz inequality and eq. 154 above, we get:
$\Big{|}\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}\rangle-\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle\Big{|}\leq\Bigg{(}\frac{C\sqrt{\frac{d+\log(2/\delta)}{n}}}{1-C\sqrt{\frac{d+\log(2/\delta)}{n}}}\Bigg{)}\|{\mathbf{x}}^{\prime}\|\|\widehat{\bm{\theta}}_{0}\|,$ (155)
with a probability of at least $1-\delta$. So if we can bound $\|\widehat{\bm{\theta}}_{0}\|$ with high probability, then we can focus on analyzing $\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle$ (which we did previously) and slightly adapting the rest of the analysis to account for this extra perturbation. We can indeed obtain sharp high-probability bounds for $\|\widehat{\bm{\theta}}_{0}\|$ by showing that it is norm-subGaussian as defined in Jin et al., (2019); we skip these calculations here as they are lengthy and tedious. Thus, as stated in Section 4.1, the crux of the proofs is in the analysis of $\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle$.
Lemma 10.
For any ${\mathbf{z}}\in\mathbb{R}^{d}$ and ${\mathbf{\Delta}}\in\mathbb{R}^{d\times d}$ such that $\|{\mathbf{\Delta}}\|<1$, we have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|\leq\Bigg{(}\frac{\|{\mathbf{\Delta}}\|}{1-\|{\mathbf{\Delta}}\|}\Bigg{)}\|{\mathbf{z}}\|.$ (156)
Proof.
We have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|$ $=\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-({\mathbf{I}}_{d}+{\mathbf{\Delta}})({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|$ (157)
$=\Big{\|}\Big{(}{\mathbf{I}}_{d}-({\mathbf{I}}_{d}+{\mathbf{\Delta}})\Big{)}({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\Big{\|}$ (158)
$\leq\|{\mathbf{\Delta}}\|\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|.$ (159)
But by the triangle inequality, we also have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|$ $\leq\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|+\|{\mathbf{z}}\|$ (160)
$\leq\|{\mathbf{\Delta}}\|\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|+\|{\mathbf{z}}\|,$ (161)
where eq. 161 follows from eq. 159. Rearranging eq. 161 yields:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|\leq\frac{\|{\mathbf{z}}\|}{1-\|{\mathbf{\Delta}}\|}.$ (162)
Plugging this into eq. 159 gives us the desired result. ∎
Appendix H Remaining Experimental Details
Here we provide the remaining details about the experiments in Section 5. Our experiments were done using TensorFlow and JAX on one 40 GB A100 GPU (per run). In all the cases, we retrain starting from random initialization rather than the previous checkpoint we converged to before RT; the former worked better than the latter. We list training details for each individual dataset next.
CIFAR-10. Optimizer is SGD with momentum $=0.9$, batch-size $=32$, number of gradient steps in each stage of training (i.e., both stages of baseline, full RT and consensus-based RT) $=21$k. We use the cosine one-cycle learning rate schedule with initial learning rate $=0.1$ for each stage of training. The number of gradient steps and initial learning rate were chosen based on the performance of the baseline method and not based on the performance of full or consensus-based RT. Standard augmentations such as random cropping, flipping and brightness/contrast change were used.
CIFAR-100. Details are the same as CIFAR-10 except that here the number of gradient steps in each stage of training $=28$k and initial learning rate $=0.005$.
AG News Subset. Small BERT model link: https://www.kaggle.com/models/tensorflow/bert/frameworks/tensorFlow2/variations/bert-en-uncased-l-4-h-512-a-8/versions/2?tfhub-redirect=true, BERT English uncased preprocessor link: https://www.kaggle.com/models/tensorflow/bert/frameworks/tensorFlow2/variations/en-uncased-preprocess/versions/3?tfhub-redirect=true. Optimizer is Adam with fixed learning rate $=1e-5$, batch size $=32$, number of epochs in each training stage $=5$.
The test accuracies without label DP for CIFAR-10, CIFAR-100 and AG News Subset are $94.13\pm 0.05\%$, $74.73\pm 0.34\%$ and $91.01\pm 0.25\%$, respectively.
Appendix I Consensus-Based Retraining Does Better than Confidence-Based Retraining
Here we compare full and consensus-based RT against another strategy for retraining which we call confidence-based retraining (RT). Specifically, we propose to retrain with the predicted labels of the samples with the top $50\%$ margin (i.e., highest predicted probability - second highest predicted probability); margin is a measure of the model’s confidence. This idea is similar to self-training’s method of sample selection in the semi-supervised setting (Amini et al.,, 2022). In Tables 7 and 8, we show results for CIFAR-10 and CIFAR-100 (in the same setting as Section 5 and Appendix H) with the smallest value of $\epsilon$ from Tables 1 and 2, respectively. Notice that consensus-based RT is clearly better than confidence-based RT.
Table 7: CIFAR-10. Test set accuracies (mean $\pm$ standard deviation). Consensus-based RT performs the best.
$\epsilon$ Baseline Full RT Consensus-based RT Confidence-based RT
1 $57.78\pm 1.13$ $60.07\pm 0.63$ $\textbf{63.84}\pm 0.56$ $62.09\pm 0.55$
Table 8: CIFAR-100. Test set accuracies (mean $\pm$ standard deviation). Again, consensus-based RT performs the best.
$\epsilon$ Baseline Full RT Consensus-based RT Confidence-based RT
3 $23.53\pm 1.01$ $24.42\pm 1.22$ $\textbf{29.98}\pm 1.11$ $24.99\pm 1.25$
Appendix J Beyond Label DP: Evaluating Retraining in the Presence of Human Annotation Errors
Even though our empirical focus in this paper has been label DP training, retraining (RT) can be employed for general problems with label noise. Here we evaluate RT in a setting with “real” label noise due to human annotation. Specifically, we focus on training a ResNet-18 model (without label DP to be clear) on the CIFAR-100N dataset introduced by Wei et al., (2021) and available on the TensorFlow website. CIFAR-100N is just CIFAR-100 labeled by humans; thus, it has real human annotation errors. The experimental setup and details are the same as CIFAR-100 (as stated in Section 5 and Appendix H); the only difference is that here we use initial learning rate $=0.01$.
In Table 9, we list the test accuracies of the baseline which is just vanilla training with the given labels, full RT and consensus-based RT, respectively. Even here with human annotation errors, consensus-based RT is beneficial.
Table 9: CIFAR-100N. Test set accuracies (mean $\pm$ standard deviation). So even with real human annotation errors, consensus-based RT improves performance.
Baseline Full RT Consensus-based RT
$55.47\pm 0.18$ $56.88\pm 0.35$ $\textbf{57.68}\pm 0.35$